### v2

# Transcendence and normality of complex numbers via Hurwitz continued fractions

Abstract.
We study the topological, dynamical, and descriptive set theoretic properties of Hurwitz continued fractions. Hurwitz continued fractions associate an infinite sequence of Gaussian integers to every complex number which is not a Gaussian rational. The resulting space of sequences of Gaussian integers $\Omega$ is not closed. By means of an algorithm, we show that $\Omega$ contains a natural subset whose closure $\overline{\mathsf{R}}$ encodes continued fraction expansions of complex numbers which are not Gaussian rationals. We prove that $(\overline{\mathsf{R}},\sigma)$ is a subshift with a feeble specification property. As an application, we determine the rank in the Borel hierarchy of the set of Hurwitz normal numbers with respect to the complex Gauss measure. We also construct a family of complex transcendental numbers with bounded partial quotients.
1. Introduction
A regular continued fraction is a finite or infinite expression of the form
(1) $a_{0}+\cfrac{1}{a_{1}+\cfrac{1}{a_{2}+\cfrac{1}{\ddots}}}$
where all the terms $a_{n}$ are integers and $a_{n}\geq 1$ for any $n\geq 1$. Regular continued fractions are a powerful tool in number theory. They provide a number system where the expansion of a real number is infinite if and only if it is irrational, and eventually periodic if and only if the number is quadratic; for example,
$\sqrt{2}=1+\cfrac{1}{2+\cfrac{1}{2+\ddots}}.$
[…]
There have been several attempts to obtain multidimensional continued fraction algorithms (see, for example, [24, 36]) but the higher dimensional theory, in comparison, is hardly as comprehensive as the one-dimensional theory. One such example is the continued fraction theory for complex plane that we focus on in this paper. In his 1887 paper [21], Hurwitz introduced a complex continued fraction algorithm where all the partial quotients $(a_{n})_{n\geq 1}$ are Gaussian integers. Despite its age, Hurwitz algorithm had not received extensive attention until recently, when interesting discoveries were made regarding its Diophantine approximation and dynamical properties [9, 12, 13, 16, 20]. Hurwitz showed that his continued fractions associate to each number $z\in\mathbb{C}\setminus\mathbb{Q}(i)$ a unique sequence $(a_{n})_{n\geq 0}$ in $\mathbb{Z}[i]$ satisfying
$z=[a_{0};a_{1},a_{2},\ldots]_{\mathbb{C}}:=\lim_{n\to\infty}a_{0}+\cfrac{1}{a_{1}+\cfrac{1}{\ddots+\cfrac{1}{a_{n}}}}.$
Hurwitz continued fractions can also be approached from a dynamical perspective. In fact, the role of the unit interval is played by the square
$\mathfrak{F}:=\left\{z\in\mathbb{C}:-\frac{1}{2}\leq\operatorname{Re}(z)<\frac{1}{2},\;-\frac{1}{2}\leq\operatorname{Im}(z)<\frac{1}{2}\right\}$
and the definition of the Hurwitz map $T$ is similar to that of $T_{\mathbb{R}}$ (see Section 3). We say a sequence $(a_{n})_{n\geq 1}$ in $\mathbb{Z}[i]$ is valid when $z=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}$ for some $z\in\mathfrak{F}\setminus\mathbb{Q}(i)$ and we denote by $\Omega$ the space of valid sequences. A central difference between regular and Hurwitz continued fractions is that, while $\mathbb{N}^{\mathbb{N}}$ is the space of valid sequences for regular continued fractions, the set $\Omega$ is not even closed in $\mathbb{Z}[i]^{\mathbb{N}}$ (see Section 4). Working with closed sets is often crucial for developing a natural dynamical perspective. One solution might involve taking the closure of $\Omega$, but this approach can result in several intricate points that prove challenging to handle. Nevertheless, it turns out that there is a naturally closed subset related to $\Omega$ that still encapsulates its fundamental characteristics. To be more precise, for each $m\in\mathbb{N}$ and $(b_{1},\ldots,b_{m})\in\mathbb{Z}[i]^{m}$, define the cylinder of level $m$ based on $\mathbf{b}$ as the set
$\mathcal{C}_{m}(\mathbf{b}):=\{z\in\mathfrak{F}:a_{1}(z)=b_{1},\ldots,a_{m}(z)=b_{m}\}$
and denote its interior by $\mathcal{C}^{\circ}_{m}(\mathbf{b})$. Let $\overline{\mathfrak{F}}$ be the closure of $\mathfrak{F}$. Define the space of sequences
$\overline{\mathsf{R}}:=\left\{(a_{n})_{n\geq 1}\in\mathbb{Z}[i]^{\mathbb{N}}:\forall n\in\mathbb{N}\quad\mathcal{C}^{\circ}_{n}(a_{1},\ldots,a_{n})\neq\varnothing\right\}$
and let $\overline{\Lambda}:\overline{\mathsf{R}}\to\overline{\mathfrak{F}}$ be the mapping acting on each $\mathbf{a}=(a_{n})_{n\geq 1}\in\overline{\mathsf{R}}$ as follows:
$\overline{\Lambda}(\mathbf{a}):=\lim_{n\to\infty}[0;a_{1},\ldots,a_{n}]_{\mathbb{C}}.$
It will be easy to show that $\overline{\Lambda}$ is well defined (see Section 4). Loosely speaking, our first result tells us that we may replace $\Omega$ with $\overline{\mathsf{R}}$.
Theorem 1.1.
$\overline{\Lambda}\left(\overline{\mathsf{R}}\right)=\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)$.
The containment $\overline{\Lambda}\left(\overline{\mathsf{R}}\right)\subseteq\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)$ will be a consequence of the approximation properties of Hurwitz continued fractions (Proposition 3.1.v). We will prove $\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)\subseteq\overline{\Lambda}\left(\overline{\mathsf{R}}\right)$, using an algorithm (Algorithm 4.3) whose input is a sequence $\mathbf{a}\in\Omega\setminus\overline{\mathsf{R}}$ and whose output is a sequence $\mathbf{b}\in\overline{\mathsf{R}}$ such that $\overline{\Lambda}(\mathbf{b})=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}$. In [9, Corollary 6.4], $\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)\subseteq\overline{\Lambda}\left(\overline{\mathsf{R}}\right)$ was obtained by different means but the argument there fails to provide for any $\mathbf{a}=(a_{n})_{n\geq 1}\in\Omega\setminus\overline{\mathsf{R}}$ a sequence $\mathbf{b}\in\overline{\mathsf{R}}$ for which $\overline{\Lambda}(\mathbf{b})=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}$.
Let $\sigma\colon\mathbb{Z}[i]^{\mathbb{N}}\to\mathbb{Z}[i]^{\mathbb{N}}$ be the left-shift map (the map that erases the first Gaussian integer). In [38], Yuri proved that $(\overline{\mathsf{R}},\sigma)$ is a sofic shift realized by a directed graph whose vertices are open prototype sets (see definition in Section 3) and with edges labelled by Gaussian integers. Yuri’s paper is concerned with the ergodic theory of a class of dynamical systems defined in compact subsets of $\mathbb{R}^{n}$, $n\in\mathbb{N}$. She considered measures equivalent to the $n$-dimensional Lebesgue measure and, thus, certain Lebesgue-null sets as well as some subsets of the associated shift space are discarded. These omissions force us to make a more detailed study of the relationship between $(\overline{\mathsf{R}},\sigma)$ and the dynamics of Hurwitz continued fraction when dealing with other problems such as those in Section 4 and Section 6.
Recall that a real number $x\in[0,1)$ is normal to base $2$ if the asymptotic frequency of every finite word in $\{0,1\}$ of length $k\in\mathbb{N}$ in the binary expansion of $x$ is exactly $2^{-k}$. In general, given a numeration system, normal numbers are those for which the asymptotic frequency of every finite word of digits is typical. For the unit interval and regular continued fractions, the notion of typical is determined by the Gauss measure of cylinders (see [3, 4, 5, 28]). In the context of Hurwitz continued fractions, we consider the unique Borel probability $T$-ergodic measure $\mu_{\operatorname{H}}$ which is equivalent to the Lebesgue measure on $\mathfrak{F}$ (see [16, 31]) and we refer to normal numbers as Hurwitz normal numbers (see Subsection 5.1 for a precise definition). As a first application of Theorem 1.1 and its proof, we determine how complicated is the set of Hurwitz normal numbers as a Borel set.
Theorem 1.2.
The set of Hurwitz normal numbers cannot be expressed as a countable intersection of open sets or as a countable union of closed (or even $G_{\delta}$) sets.
In fact, we will establish a more robust result by categorizing the set of normal numbers within the framework of the Borel hierarchy. The Borel hierarchy serves as a natural rank, quantifying the number of quantifiers required to construct a Borel set, commencing from open sets. In some sense, this measures how difficult it is to determine whether a complex number is a Hurwitz normal number. A related result was obtained by Dani and Nogueira when they were studying the image of $\mathbb{Z}[i]^{2}$ under binary quadratic forms [13, Propositions 8.5, 8.6].
Theorem 1.2 is a consequence of the dynamical structure of $\overline{\mathsf{R}}$ with respect to the left-shift map. We prove a weak form of specification introduced by Airey, Jackson, Kwietniak, and Mance in [4].
Theorem 1.3.
The subshift $(\overline{\mathsf{R}},\sigma)$ has the right feeble-specification property.
The subshift $(\overline{\mathsf{R}},\sigma)$ is not a full shift, in other words, $\overline{\mathsf{R}}$ is not of the form $\mathscr{A}^{\mathbb{N}}$ for some $\mathscr{A}\subseteq\mathbb{Z}[i]$. Nonetheless, Theorem 1.3 shows that the system $(\overline{\mathsf{R}},\sigma)$ still has a considerable level of freedom in terms of specification. The specification property was introduced by Bowen in his seminal paper [6]. Bowen’s Specification Theorem asserts that all transitive hyperbolic systems possess the “specification property”, wherein complete independence is established among finite subsets with bounded gaps. The specification and its variations have proven to be potent tools for unlocking access to various other properties, such as the uniqueness of measures of maximal entropy [6, 11], universality [32], and the intricate structure of generic points [4, 14].
The second application of Theorem 1.1 is related to transcendental numbers. In [2], Adamczewski and Bugeaud solved Cobham’s Conjecture when they proved that, for any $b\in\mathbb{N}_{\geq 2}$, a real number whose expansion in base $b$ is non-periodic and generated by finite automaton must be transcendental. They extended their work to continued fractions [1] and, afterwards in [8], Bugeaud gave a combinatorial condition on non-periodic continued fraction ensuring its transcendence. This condition can be expressed in an extremely neat manner using the repetition exponent introduced by Bugeaud and Kim [10]. Given a finite alphabet $\mathscr{A}$ and a sequence $\mathbf{a}=(a_{n})_{n\geq 1}$ in $\mathscr{A}$, for each $n\in\mathbb{N}$ define
$r(n,\mathbf{a}):=\min\left\{m\in\mathbb{N}:\exists j\in\{1,\ldots,m-n\}\quad(a_{j},\ldots,a_{j+n-1})=(a_{m},\ldots,a_{m+n-1})\right\}.$
The repetition exponent of $\mathbf{a}$ is
$\operatorname{rep}(\mathbf{a}):=\liminf_{n\to\infty}\frac{r(n,\mathbf{a})}{n}.$
Bugeaud proved that for any bounded non-periodic sequence $\mathbf{a}=(a_{n})_{n\geq 1}\in\mathbb{N}^{\mathbb{N}}$, if $\operatorname{rep}(\mathbf{a})<\infty$, the continued fraction (1) is transcendental for any $a_{0}\in\mathbb{Z}$ [8, Theorem 1.3]. In [19], the second author obtained a complex version of this result under the addition assumption $\min_{n}|a_{n}|\geq\sqrt{8}$. In this paper, we apply the algorithm from the proof of Theorem 1.1 to construct a family of transcendental numbers whose partial quotients $(a_{n})_{n\geq 1}$ are bounded and such that $|a_{n}|=2$ for infinitely many $n\in\mathbb{N}$. Schmidt’s Subspace Theorem, a deep result in Diophantine approximation, lies at the heart of this construction.
Theorem 1.4.
Let $\mathbf{B}:=(B_{n})_{n\geq 1}$ be a bounded sequence in $\mathbb{Z}$ which is not periodic and such that
$\min_{n\in\mathbb{N}}|B_{n}|\geq 3\quad\text{ and }\quad\operatorname{rep}(\mathbf{B})<\infty.$
Then, $\zeta=[0;-2,1+iB_{1},-2,1+iB_{2},-2,1+iB_{3},-2,1+iB_{4},\ldots]_{\mathbb{C}}$ is transcendental.
To sum up, these are our applications:
  1. (1)
    The set $\overline{\mathsf{R}}:=\left\{(a_{n})_{n\geq 1}\in\mathbb{Z}[i]^{\mathbb{N}}:\forall n\in\mathbb{N}\quad\mathcal{C}^{\circ}_{n}(a_{1},\ldots,a_{n})\neq\varnothing\right\}$ is closed and dynamically chaotic in the sense that it has a weak type of specification.
  2. (2)
    The set of Hurwitz normal numbers is classified in the third level of the Borel hierarchy.
  3. (3)
    We construct a family of transcendental complex numbers with bounded partial quotients.
The structure of the paper is as follows. First, in Section 2, we establish notation and conventions. In Section 3, we introduce the basic definitions of Hurwitz continued fractions, we discuss the classification of cylinders and prototype sets, and we give some basic results. In Section 4 we prove Theorem 1.1 and state Algorithm 4.3 the algorithm without proof. Section 5 starts with some basic notions on Borel Hierarchy and the feeble-specification property. Later, we show Theorem 1.3 and then Theorem 5.1 (which implies Theorem 1.2). In Section 6 we show Theorem 1.4. Lastly, Section 7 contains the proof Algorithm 4.3.
2. Notation and conventions
We gather some frequently used notations.
  1. i.
    By natural numbers $\mathbb{N}$ we mean the set of positive integers and we write $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$.
  2. ii.
    The real part of a complex number $z$ is denoted by $\operatorname{Re}(z)$, its imaginary part by $\operatorname{Im}(z)$, its complex conjugate by $\overline{z}$, and $\operatorname{Pm}(z):=\min\{|\operatorname{Re}(z)|,|\operatorname{Im}(z)|\}$.
  3. iii.
    If $X,Y$ are sets, $A\subseteq X$, $B\subseteq Y$ and $f:X\to Y$ is any function, we define
    $f[A]:=\{f(a):a\in A\}\quad\text{and}\quad f^{-1}[B]:=\{a\in A:f(a)\in B\}.$
  4. iv.
    $\mathbb{Z}[i]:=\{a+ib\in\mathbb{C}:a,b\in\mathbb{Z}\}$, and $\mathbb{Q}(i):=\{a+ib\in\mathbb{C}:a,b\in\mathbb{Z}\in\mathbb{Q}\}$.
  5. v.
    For any $a\in\mathbb{Z}[i]$, the function $\tau_{a}:\mathbb{C}\to\mathbb{C}$ is given by $\tau_{a}(z)=z+a$, $z\in\mathbb{C}$. By $\iota$ we mean the complex inversion: $\iota(z)=z^{-1}$ for $z\in\mathbb{C}\setminus\{0\}$.
  6. vi.
    Given finitely many functions $h_{1},\ldots,h_{n}$ that are either $\iota$, $\tau_{a}$ for some $a\in\mathbb{Z}[i]$ or some element in $\mathsf{Di}_{8}$ (see Section 3), we write $h_{n}\cdots h_{1}:=h_{n}\circ\cdots\circ h_{1}$.
  7. vii.
    For $z\in\mathbb{C}$ and $\rho>0$, we write
    $\mathbb{D}(z;\rho):=\{w\in\mathbb{C}:|z-w|<\rho\},\quad$ $\quad\mathbb{D}(z):=\mathbb{D}(z;1),$
    $\overline{\mathbb{D}}(z;\rho):=\{w\in\mathbb{C}:|z-w|\leq\rho\},\quad$ $\quad\overline{\mathbb{D}}(z):=\overline{\mathbb{D}}(z;1),$
    $C(z;\rho):=\{w\in\mathbb{C}:|z-w|=\rho\},\quad$ $\quad C(z):=C(z;1).$
  8. viii.
    If $X$ is a topological space and $A\subseteq X$, we let $\operatorname{Cl}_{X}(A)$ (resp. $\operatorname{Int}_{X}(A)$ ) be the closure of $A$ (resp. the interior). We omit the sub-index $X$ if the space is clear.
  9. ix.
    The term word typically refers to a finite set of ordered symbols, while a sequence denotes an infinite one. However, in certain situations where the set may be finite or infinite, we use both terms interchangeably.
  10. x.
    We denote the empty word by $\epsilon$.
  11. xi.
    If $\mathscr{A}$ is a non-empty set, $m\in\mathbb{N}$, and $\mathbf{a}=(a_{1},\ldots,a_{m})\in\mathscr{A}^{m}$, the length of $\mathbf{a}$ is $|\mathbf{a}|:=m$. We further define $|\epsilon|=0$. If $\mathbf{b}=(b_{n})_{n\geq 1}$ is an infinite or finite word in $\mathscr{A}$, we write
    $\mathbf{a}\mathbf{b}:=(a_{1},\ldots,a_{m},b_{1},b_{2},\ldots).$
  12. xii.
    If $\mathscr{A}$ is a non-empty set and $\mathbf{a}=(a_{n})_{n\geq 1}\in\mathscr{A}^{\mathbb{N}}$, a factor of $\mathbf{a}$ is a subword that appears in any position and a prefix is a factor that appears at the beginning of $\mathbf{a}$. For each $m\in\mathbb{N}$ the prefix of length $m$ of $\mathbf{a}$ is $\operatorname{pref}(\mathbf{a};m):=(a_{1},\ldots,a_{m})$.
We use implicitly the formulas for the inversion of discs and lines in the complex plane throughout the paper. Namely, for $k\in\mathbb{R}$, we have
(2) $\iota\left[\{z\in\mathbb{C}:\operatorname{Re}(z)=k\}\right]=C\left(\frac{1}{2k};\frac{1}{2|k|}\right),\quad\iota\left[\{z\in\mathbb{C}:\operatorname{Im}(z)=k\}\right]=C\left(\frac{-i}{2k};\frac{1}{2|k|}\right)$
Also, if $z_{0}\in\mathbb{C}$ and $\rho>0$ are such that $\rho\neq|z_{0}|$ then
(3) $\iota\left[C(z_{0};\rho)\right]=C\left(\frac{\overline{z_{0}}}{|z|^{2}-\rho^{2}};\frac{\rho}{|\rho^{2}-|z_{0}|^{2}|}\right).$
3. Hurwitz continued fractions
We start this section by defining Hurwitz continued fractions, the oldest complex continued fraction algorithm; other algorithms have been discussed in [13, 18, 30, 34]. Let $\lfloor\,\cdot\,\rfloor:\mathbb{R}\to\mathbb{Z}$ be the usual floor function. For $z\in\mathbb{C}$, the nearest Gaussian integer $[z]$ to $z$ is
$[z]:=\left\lfloor\operatorname{Re}(z)+\frac{1}{2}\right\rfloor+i\left\lfloor\operatorname{Im}(z)+\frac{1}{2}\right\rfloor.$
We define
$\mathfrak{F}=\{z\in\mathbb{C}:[z]=0\}.$
The Hurwitz map, $T:\mathfrak{F}\to\mathfrak{F}$, is given by
$\quad T(z)=\begin{cases}z^{-1}-[z^{-1}],&\text{ if }z\neq 0,\\ 0,&\text{ if }z=0.\end{cases}$
For any $z\in\mathfrak{F}\setminus\{0\}$, we define $a_{0}(z)=0$, $a_{1}(z)\colon=[z^{-1}]$ and, if $T^{n-1}(z)\neq 0$, we put $a_{n}(z)\colon=a_{1}(T^{n-1}(z))$ for $n\in\mathbb{N}$. As it is customary, the exponent in $T$ denotes iteration. For a general $z\in\mathbb{C}$, we define $a_{0}(z)=[z]$ and $a_{n}(z)=a_{n}(z-[z])$, $n\in\mathbb{N}$. The Hurwitz continued fraction (HCF) of $z$ is
$[a_{0}(z);a_{1}(z),a_{2}(z),\ldots]_{\mathbb{C}}\colon=\cfrac{1}{a_{1}(z)+\cfrac{1}{a_{2}(z)+\cfrac{1}{\ddots}}}.$
We will restrict our attention to $z\in\mathfrak{F}$. Note that
(4) $[0;a_{1}(T(z)),a_{2}(T(z)),\ldots]_{\mathbb{C}}=[0;a_{2}(z),a_{3}(z)\ldots]_{\mathbb{C}}.$
The terms of the sequence $(a_{n}(z))_{n\geq 1}$ are the partial quotients of $z$. Define the sequences of Gaussian integers $(p_{n})_{n\geq 0}$, $(q_{n})_{n\geq 0}$ by:
(5) $\begin{pmatrix}p_{-1}&p_{-2}\\ q_{-1}&q_{-2}\end{pmatrix}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\qquad\text{ and }\qquad\forall n\in\mathbb{N}_{0}\quad\begin{pmatrix}p_{n}\\ q_{n}\end{pmatrix}=\begin{pmatrix}p_{n-1}&p_{n-2}\\ q_{n-1}&q_{n-2}\end{pmatrix}\begin{pmatrix}a_{n}\\ 1\end{pmatrix}.$
The general theory of continued fractions (see, for example, [27, Section 2]) tells us that
$\forall n\in\mathbb{N}\quad\frac{p_{n}}{q_{n}}=[0;a_{1}(z),a_{2}(z),\ldots,a_{n}(z)]_{\mathbb{C}}:=\cfrac{1}{a_{1}(z)+\cfrac{1}{a_{2}(z)+\cfrac{1}{\ddots+\cfrac{1}{a_{n}(z)}}}}.$
The quotients $\frac{p_{n}}{q_{n}}$ are the convergents of $z$. From (5) we get
(6) $\forall n\in\mathbb{N}\quad q_{n}p_{n-1}-q_{n-1}p_{n}=(-1)^{n}.$
There are complex numbers $\xi$ for which there are more than one Gaussian integer $a_{0}$ solving
(7) $|\xi-a_{0}|=\min\left\{|\xi-a|:a\in\mathbb{Z}[i]\right\}.$
For such $\xi$, we have defined $[\xi]$ to be the solution of (7) with the largest real and imaginary parts. In many situations, the breaking ties criterion is irrelevant for it applies on a set of Hausdorff dimension $1$, hence Lebesgue null. However, it ensures the uniqueness of the HCF expansion. If we are willing to loosen slightly the uniqueness of the expansion, Theorem 1.1 and its proof imply that it makes no difference what solution of (7) we choose.
We will use continued fractions which are not necessarily HCFs and denote them by
$[b_{0};b_{1},b_{2},\ldots]:=b_{0}+\cfrac{1}{b_{1}+\cfrac{1}{b_{2}+\ddots}}.$
We use the notation $[b_{0};b_{1},b_{2},\ldots]_{\mathbb{C}}$ for HCFs.
Next, we recall some elementary facts about HCFs. Define $\mathscr{D}:=\mathbb{Z}[i]\setminus\{0,1,i,-1,-i\}$.
Proposition 3.1.
Let $z=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}\in\mathfrak{F}\setminus\mathbb{Q}(i)$ and let $(p_{n})_{n\geq 0}$, $(q_{n})_{n\geq 0}$ be as in (5).
  1. i.
    [21, Section I] For every $n$, we have $a_{n}\in\mathscr{D}$.
  2. ii.
    [21, Section I] The sequence $(|q_{n}|)_{n\geq 0}$ is strictly increasing.
  3. iii.
    [13, Corollary 5.3] If $\psi\colon=\sqrt{\frac{1+\sqrt{5}}{2}}$, then $|q_{n}|\geq\psi^{n-1}$ for all $n\in\mathbb{N}$.
  4. iv.
    [13, Corollary 3.3] We have
    $\forall n\in\mathbb{N}\quad z=\frac{(a_{n+1}+[0;a_{n+2},a_{n+3},\ldots])p_{n}+p_{n-1}}{(a_{n+1}+[0;a_{n+2},a_{n+3},\ldots])q_{n}+q_{n-1}}.$
  5. v.
    [29, Theorem 1] If $z=[0;a_{1},a_{2},\ldots]\in\mathfrak{F}\setminus\mathbb{Q}(i)$, then
    $\forall n\in\mathbb{N}\quad\left|z-\frac{p_{n}}{q_{n}}\right|<\frac{1}{|q_{n}|^{2}}.$
  6. vi.
    [19, Lemma 4.4] There is a universal constant $\lambda>0$ such that
    $\forall n\in\mathbb{N}\quad\left|z-\frac{p_{n}}{q_{n}}\right|<\frac{\lambda}{|q_{n}q_{n+1}|^{2}}.$
  7. vii.
    [20, Lemma 2.1.(g)] There are constants $c_{1},c_{2}>0$ such that for all $n,k\in\mathbb{N}$ we have
    $c_{1}\leq\frac{|q_{n+k}(a_{1},\ldots,a_{n+k})|}{|q_{n}(a_{1},\ldots,a_{n})||q_{k}(a_{n+1},\ldots,a_{n+k})|}\leq c_{2}.$
3.1. HCFs Cylinders
Cylinders and prototype sets are the building blocks of the geometric theory of HCFs. Given $n\in\mathbb{N}$ and $\mathbf{a}=(a_{1},\ldots,a_{n})\in\mathscr{D}^{n}$, the prototype set $\mathfrak{F}_{n}(\mathbf{a})$ is defined by
$\mathfrak{F}_{n}(\mathbf{a})\colon=T^{n}[\mathcal{C}_{n}(\mathbf{a})].$
We denote the interior (resp. closure) of $\mathcal{C}_{n}(\mathbf{a})$ by $\mathcal{C}^{\circ}_{n}(\mathbf{a})$ (resp. $\overline{\mathcal{C}}_{n}(\mathbf{a})$) and similar conventions apply on $\mathfrak{F}$. For brevity, by open cylinders we mean the interior of a regular cylinder and similarly for prototype sets. Cylinders induce a natural classification of finite words in $\mathscr{D}$. We say that $\mathbf{a}$ is
  1. i.
    valid if $\mathcal{C}_{n}(\mathbf{a})\neq\varnothing$;
  2. ii.
    regular if $\mathcal{C}^{\circ}_{n}(\mathbf{a})\neq\varnothing$;
  3. iii.
    irregular if $\mathcal{C}_{n}(\mathbf{a})\neq\varnothing$ and $\mathcal{C}^{\circ}_{n}(\mathbf{a})=\varnothing$;
  4. iv.
    extremely irregular if #$\#\mathcal{C}_{n}(\mathbf{a})=1$.
Naturally, #$\#A$ stands for the number of elements in a set $A$. We denote by $\Omega(n)$ (resp. $\mathsf{R}(n)$, $\mathsf{Ir}(n)$, $\mathsf{EI}(n)$) the set of valid (resp. regular, irregular, extremely irregular) words of length $n$. We extend the notions regular, irregular, and extremely irregular to cylinders and prototype sets in an obvious way. It is well known that there are only finitely many prototype sets (see, for example, [16]). Actually, an open prototype set is one of the following sets or one of their right-angled rotations (see Figure 1):
$\mathfrak{F}^{\circ},\quad\mathfrak{F}^{\circ}_{1}(-2)=\mathfrak{F}^{\circ}\setminus\overline{\mathbb{D}}(1),$
$\mathfrak{F}^{\circ}_{1}(-2+i)=\mathfrak{F}^{\circ}\setminus\overline{\mathbb{D}}(1-i),\quad\mathfrak{F}^{\circ}_{1}(-1+i)=\mathfrak{F}^{\circ}\setminus\left(\overline{\mathbb{D}}(1)\cup\overline{\mathbb{D}}(-i)\right).$
Figure 1. From left to right: $\mathfrak{F}^{\circ}$, $\mathfrak{F}^{\circ}_{1}(-2)$, $\mathfrak{F}^{\circ}_{1}(-2+i)$, $\mathfrak{F}^{\circ}_{1}(-1+i)$.
A sequence $\mathbf{a}=(a_{n})_{n\geq 1}\in\mathscr{D}^{\mathbb{N}}$ is
  1. i.
    valid if it is the HCF of some $z\in\mathfrak{F}$;
  2. ii.
    regular if it is valid and $(a_{1},\ldots,a_{n})\in\mathsf{R}(n)$ for all $n\in\mathbb{N}$;
  3. iii.
    irregular if it is valid and $(a_{1},\ldots,a_{n})\in\mathsf{Ir}(n)$ for some $n\in\mathbb{N}$;
  4. iv.
    extremely irregular if it is valid and $(a_{1},\ldots,a_{n})\in\mathsf{EI}(n)$ for some $n\in\mathbb{N}$.
We denote by $\Omega$ (resp. $\mathsf{R}$, $\mathsf{Ir}$, $\mathsf{EI}$) the set of valid (resp. regular, irregular, extremely irregular) sequences.
A significant part of our calculations and our arguments is based on an observation stemming from the definitions of cylinders: if $(a_{n})_{n\geq 1}$ is a sequence in $\mathscr{D}$, then
$\mathcal{C}_{1}(a_{1})=\mathfrak{F}\cap\iota\tau_{a_{1}}[\mathfrak{F}],\quad\mathcal{C}_{2}(a_{1},a_{2})=\mathfrak{F}\cap\iota\tau_{a_{1}}[\mathfrak{F}]\cap\iota\tau_{a_{2}}\iota\tau_{a_{1}}[\mathfrak{F}];$
in general,
(8) $\forall n\in\mathbb{N}\quad\mathcal{C}_{n}(a_{1},\ldots,a_{n})=\mathfrak{F}\cap\iota\tau_{a_{1}}[\mathfrak{F}]\cap\ldots\cap(\iota\tau_{a_{1}}\cdots\iota\tau_{a_{n}})[\mathfrak{F}].$
Since $\iota$ and $\tau_{a}$ are homeomorphisms from the Riemann sphere onto itself, we may replace $\mathfrak{F}$ and $\mathcal{C}$ with $\mathfrak{F}^{\circ}$ and $\mathcal{C}^{\circ}$, respectively, in (8). It follows from (8) that every factor of a regular word (finite or infinite) is again regular. We also deduce the next proposition.
Proposition 3.2.
Put $\mathfrak{F}_{0}(\epsilon):=\mathfrak{F}$. Every sequence $(a_{n})_{n\geq 1}$ in $\mathscr{D}$ satisfies
$\forall n\in\mathbb{N}\quad\mathfrak{F}_{n}(a_{1},\ldots,a_{n})=\tau_{-a_{n}}\iota\left[\mathfrak{F}_{n-1}(a_{1},\ldots,a_{n-1})\cap\mathcal{C}_{1}(a_{n})\right].$
3.2. Regular cylinders
We recall two important properties of regular cylinders. The first property tells us that every infinite sequence in $\mathscr{D}$ with large terms is regular. The second proposition establishes symmetries of the family of regular cylinders.
Proposition 3.3 ([9, Proposition 4.3]).
If $\mathbf{a}=(a_{n})_{n\geq 1}$ is such that $\min_{n\in\mathbb{N}}|a_{n}|\geq\sqrt{8}$, then $\mathbf{a}\in\mathsf{R}$ and $\mathfrak{F}_{n}(a_{1},\ldots,a_{n})=\mathfrak{F}$ for all $n\in\mathbb{N}$.
Let $\mathsf{rot},\mathsf{mir}_{1}:\mathbb{C}\to\mathbb{C}$ be given by
$\forall z\in\mathbb{C}\quad\mathsf{rot}(z):=iz\quad\text{ and }\quad\mathsf{mir}_{1}(z):=\overline{z}$
and call $\mathsf{Di}_{8}$ the group of isometries they generate. Obviously, $\mathsf{Di}_{8}$ is the dihedral group of order $8$. Proposition 4.5 and Corollary 4.6 in [9] say that any $f\in\mathsf{Di}_{8}$ maps open cylinders onto open cylinders of the same level. These results are proven by explicitly determining the image of an open cylinder under $\mathsf{rot}$ and $\mathsf{mir}_{1}$. In particular, if $\mathsf{mir}_{2}:=\mathsf{rot}^{2}\mathsf{mir}_{1}$, $z\mapsto-\overline{z}$, and $\mathsf{Mir}\subseteq\mathsf{Di}_{8}$ is the subgroup generated by $\mathsf{mir}_{1}$ and $\mathsf{mir}_{2}$, we have the next result.
Proposition 3.4.
If $n\in\mathbb{N}$, $\mathbf{a}=(a_{1},\ldots,a_{n})\in\mathsf{R}(n)$, and $\mathsf{mir}\in\mathsf{Mir}$, then
$\mathsf{mir}\left[\mathcal{C}^{\circ}_{n}(\mathbf{a})\right]=\mathcal{C}^{\circ}_{n}(\mathsf{mir}(a_{1}),\ldots,\mathsf{mir}(a_{n})).$
3.3. Irregular cylinders
For any pair of complex numbers $z$ and $w$, write
$[z,w):=\{z+t(w-z):t\in[0,1)\}.$
Irregular cylinders, and thus prototype sets, exist. Certainly, if
$\alpha\colon=\frac{2-\sqrt{3}}{2}\approx 0.13397,$
for every $m\in\mathbb{Z}$ with $|m|\geq 2$ we have
$\mathfrak{F}_{2}(-2,1+mi)=\begin{cases}\left[-\frac{1}{2}-i\alpha,-\frac{1}{2}+\frac{i}{2}\right),\ \text{\rm if }\ m=2,\\[8.61108pt] \left[-\frac{1}{2}-\frac{i}{2},-\frac{1}{2}+i\alpha\right),\ \text{\rm if }\ m=-2,\\[8.61108pt] \left[-\frac{1}{2}-\frac{i}{2},-\frac{1}{2}+\frac{i}{2}\right),\ \text{\rm if }\ |m|\geq 3\end{cases}$
(see [9, Section 4]). Additionally, if
$\zeta_{1}\colon=-\frac{1}{2}+i\alpha,\quad\zeta_{2}\colon=-\frac{1}{2}-i\alpha,\quad\zeta_{3}\colon=-\alpha-\frac{i}{2},\quad\zeta_{4}\colon=\alpha-\frac{i}{2},$
straightforward calculations yield $a_{1}(\zeta_{1})=-2$, $a_{1}(\zeta_{2})=-2+i$, $a_{1}(\zeta_{3})=2i$, $a_{1}(\zeta_{4})=1+2i$ and
(9) $T(\zeta_{1})=\zeta_{4},\quad T(\zeta_{2})=\zeta_{4},\quad T(\zeta_{3})=\zeta_{2},\quad T(\zeta_{4})=\zeta_{2}.$
As a consequence, we have the HCF expansions
$\zeta_{1}$ $=[0;-2,1+2i,-2+i,1+2i,-2+i,1+2i,-2+i,\ldots]_{\mathbb{C}},$
$\zeta_{2}$ $=[0;-2+i,1+2i,-2+i,1+2i,-2+i,1+2i,-2+i,\ldots]_{\mathbb{C}},$
$\zeta_{3}$ $=[0;2i,-2+i,1+2i,-2+i,1+2i,-2+i,1+2i,\ldots]_{\mathbb{C}},$
$\zeta_{4}$ $=[0;1+2i,-2+i,1+2i,-2+i,1+2i,-2+i,1+2i,\ldots]_{\mathbb{C}}.$
Using (3) and (8), we may show that the sequences $(a_{n}(\zeta_{j}))_{n\geq 1}$, $j\in\{1,2,3,4\}$, are extremely irregular. In fact, by carefully tracking the image under $T$ of the boundaries of prototype sets, we may even show that the HCF of a complex number $\xi\in\mathfrak{F}$ is extremely irregular if and only if $T^{n}(\xi)=\zeta_{4}$ for some $n\in\mathbb{N}$. Figure 2 depicts the partition of $\mathfrak{F}$ induced by the cylinders $\{\mathcal{C}_{1}(a):a\in\mathscr{D}\}$ and the location of $\zeta_{1},\zeta_{2},\zeta_{3},\zeta_{4}$ in $\mathfrak{F}$.
Figure 2. The cylinders $\mathcal{C}_{1}(a)$, $a\in\mathscr{D}$, and the numbers $\zeta_{1},\zeta_{2},\zeta_{3},\zeta_{4}$.
Lemma 3.5.
Let $n\in\mathbb{N}$ be arbitrary. If $\mathbf{a}\in\mathsf{R}(n)$ and $b\in\mathscr{D}$ are such that $\mathbf{a}b\in\mathsf{Ir}(n+1)$, then $\mathfrak{F}^{\circ}_{n}(\mathbf{a})$ is one of the next sets:
$\mathfrak{F}^{\circ}\setminus\overline{\mathbb{D}}(1),\quad\mathfrak{F}^{\circ}\setminus\overline{\mathbb{D}}(-i)$ $,\quad\mathfrak{F}^{\circ}\setminus\left(\overline{\mathbb{D}}(1)\cup\overline{\mathbb{D}}(-i)\right),$
$\mathfrak{F}^{\circ}\setminus\left(\overline{\mathbb{D}}(-1)\cup\overline{\mathbb{D}}(-i)\right),$ $\quad\mathfrak{F}^{\circ}\setminus\left(\overline{\mathbb{D}}(1)\cup\overline{\mathbb{D}}(i)\right).$
Proof.
The proof amounts to check that every unmentioned option for $\mathfrak{F}^{\circ}_{n}(\mathbf{a})\neq\varnothing$ satisfies
$\mathbf{a}b\in\Omega(n+1)\quad\text{ implies }\quad\mathbf{a}b\in\mathsf{R}(n+1),$
which follows from the stronger statement
$\iota\left[\overline{\mathfrak{F}}_{n}(\mathbf{a})\right]\cap\tau_{b}\left[\mathfrak{F}\right]\neq\varnothing\quad\text{ implies }\quad\iota\left[\mathfrak{F}^{\circ}_{n}(\mathbf{a})\right]\cap\tau_{b}\left[\mathfrak{F}^{\circ}\right]\neq\varnothing.$
We rely on pictures in order to avoid long yet elementary computations.
  1. i.
    If $\mathfrak{F}^{\circ}_{n}(\mathbf{a})=\mathfrak{F}^{\circ}$, then $\iota[\mathfrak{F}^{\circ}_{n}(\mathbf{a})]\cap\tau_{b}[\mathfrak{F}^{\circ}]\neq\varnothing$ for all $b\in\mathscr{D}$ (see Figure 3).
  2. ii.
    When $\mathfrak{F}^{\circ}_{n}(\mathbf{a})=\mathfrak{F}^{\circ}\setminus\overline{\mathbb{D}}(1+i)$ we have $\iota[\overline{\mathfrak{F}}_{n}(\mathbf{a})]=\iota[\overline{\mathfrak{F}}]\setminus\mathbb{D}(1-i)$. We conclude from Figure 4 that $\iota\left[\overline{\mathfrak{F}}_{n}(\mathbf{a})\right]\cap\tau_{-b}\iota[\mathfrak{F}]\neq\varnothing$ if and only if $b\neq 1-i$, in which case $\iota\left[\mathfrak{F}^{\circ}_{n}(\mathbf{a})\right]\cap\tau_{-b}\iota[\mathfrak{F}^{\circ}]\neq\varnothing$. A similar argument holds when $\mathfrak{F}^{\circ}_{n}(\mathbf{a})=\mathfrak{F}^{\circ}\setminus i^{k}\overline{\mathbb{D}}(1+i)$ for some $k\in\{2,3,4\}$.
  3. iii.
    If $\mathfrak{F}^{\circ}_{n}(\mathbf{a})=\mathfrak{F}^{\circ}\setminus\overline{\mathbb{D}}(-1)$, then
    $\iota\left[\overline{\mathfrak{F}}_{n}(\mathbf{a})\right]=\iota\left[\overline{\mathfrak{F}}\right]\cap\left\{z\in\mathbb{C}:\operatorname{Re}(z)\geq-\frac{1}{2}\right\}$
    and, by Figure 5, $\iota\left[\overline{\mathfrak{F}}_{n}(\mathbf{a})\right]\neq\varnothing$ if and only if $\operatorname{Re}(b)\geq 0$. Figure 5 also shows that $\operatorname{Re}(b)\geq 0$ yields $\iota[\mathfrak{F}^{\circ}_{n}(\mathbf{a})]\cap\tau_{b}[\mathfrak{F}^{\circ}]\neq\varnothing$. A similar argument holds when $\mathfrak{F}^{\circ}_{n}(\mathbf{a})=\mathfrak{F}^{\circ}\setminus\overline{\mathbb{D}}(i)$.
  4. iv.
    When $\mathfrak{F}^{\circ}_{n}(\mathbf{a})=\mathfrak{F}^{\circ}\setminus(\overline{\mathbb{D}}(i)\cup\overline{\mathbb{D}}(-1))$, by Figure 6 we have $\tau_{b}\left[\mathfrak{F}^{\circ}_{n}(\mathbf{a})\right]\cap\iota[\overline{\mathfrak{F}}]\neq\varnothing$ if and only if $\operatorname{Re}(b)\geq 0$ and $\operatorname{Im}(b)\geq 0$. For such $b$ we have $\iota[\mathfrak{F}^{\circ}_{n}(\mathbf{a})]\cap\tau_{b}[\mathfrak{F}^{\circ}]\neq\varnothing$.
This shows that if $\mathbf{a}\in\mathsf{R}(n)$ and $\mathfrak{F}^{\circ}_{n}(\mathbf{a})\neq\varnothing$ is not one of the sets in the statement, then $\mathbf{a}b\in\Omega(n+1)$ implies $\mathbf{a}b\in\mathsf{R}(n+1)$. ∎
Figure 3. The set $\iota[\mathfrak{F}^{\circ}]$.
Refer to captionRefer to caption
Figure 4. The sets $\mathfrak{F}^{\circ}\setminus\overline{\mathbb{D}}(1+i)$ (left) and $\iota[\overline{\mathfrak{F}}\setminus\mathbb{D}(1+i)]$ (right).
Refer to captionRefer to caption
Figure 5. The sets $\mathfrak{F}^{\circ}\setminus\overline{\mathbb{D}}(-1)$ (left) and $\iota[\overline{\mathfrak{F}}\setminus\mathbb{D}(-1)]$ (right).
Refer to captionRefer to caption
Figure 6. The sets $\mathfrak{F}^{\circ}\setminus(\overline{\mathbb{D}}(i)\cup\overline{\mathbb{D}}(-1))$ (left) and $\iota[(\overline{\mathfrak{F}}\setminus(\mathbb{D}(i)\cup\mathbb{D}(-1)))]$ (right).
4. Proof of Theorem 1.1
4.1. The shift space
Endow $\mathscr{D}^{\mathbb{N}}$ with the product topology assuming each factor has a discrete topology. Then, $\mathscr{D}^{\mathbb{N}}$ can be metricized with the metric $\operatorname{dist}$ given by
$\forall\mathbf{a},\mathbf{b}\in\mathscr{D}^{\mathbb{N}}\quad\operatorname{dist}(\mathbf{a},\mathbf{b}):=\begin{cases}0,\text{ if }\mathbf{a}=\mathbf{b},\\ 2^{-\min\{n\in\mathbb{N}:\operatorname{pref}(\mathbf{a};n)\neq\operatorname{pref}(\mathbf{b};n)\}},\text{ if }\mathbf{a}\neq\mathbf{b}.\end{cases}$
Recall that $\sigma\colon\mathscr{D}^{\mathbb{N}}\to\mathscr{D}^{\mathbb{N}}$ is the map associating to each $\mathbf{x}=(x_{n})_{n\geq 1}\in\mathscr{D}^{\mathbb{N}}$ the sequence $\sigma(\mathbf{x})=(x_{j})_{j\geq 2}$. We link the dynamical systems $(\Omega,\sigma)$ and $(\mathfrak{F}\setminus\mathbb{Q}(i),T)$ through the map $\Lambda:\Omega\to\mathfrak{F}\setminus\mathbb{Q}(i)$ given by
$\forall\mathbf{a}=(a_{n})_{n\geq 1}\in\Omega\quad\Lambda(\mathbf{a}):=[0;a_{1},a_{2},a_{3},\ldots]_{\mathbb{C}}.$
Let $\mathbf{a}\in\Omega$ and $\varepsilon>0$ be arbitrary. Take $m\in\mathbb{N}$ such that, for $\psi$ as in Proposition 3.1.iii, $\frac{2}{\psi^{m-1}}<\varepsilon$ and denote by $p_{m}/q_{m}$ the $m$-th convergent of $\Lambda(\mathbf{a})$. If $\mathbf{b}\in\Omega$ is such that $\operatorname{dist}(\mathbf{a},\mathbf{b})<2^{-m}$, then $\operatorname{pref}(\mathbf{a};m)=\operatorname{pref}(\mathbf{b};m)$ and, by Proposition 3.1.v,
$\left|\Lambda(\mathbf{a})-\Lambda(\mathbf{b})\right|\leq\left|\Lambda(\mathbf{a})-\frac{p_{m}}{q_{m}}\right|+\left|\Lambda(\mathbf{b})-\frac{p_{m}}{q_{m}}\right|\leq\frac{2}{|q_{m}|^{2}}\leq\frac{2}{\psi^{m-1}}<\varepsilon.$
In other words, $\Lambda$ is uniformly continuous. Call $\Lambda|_{\mathsf{R}}$ the restriction of $\Lambda$ to $\mathsf{R}$ and let $\overline{\Lambda}$ be the unique continuous extension of $\Lambda|_{\mathsf{R}}$ to $\overline{\mathsf{R}}:=\operatorname{Cl}_{\mathscr{D}^{\mathbb{N}}}(\mathsf{R})$. We may now state a lemma containing the core of Theorem 1.1. We will prove it by constructing algorithmically the sequence $\mathbf{b}$ from $\mathbf{a}$.
Lemma 4.1.
If $\mathbf{a}\in\mathsf{Ir}$, then there is some $\mathbf{b}\in\overline{\mathsf{R}}$ such that $\overline{\Lambda}(\mathbf{b})=\Lambda(\mathbf{a})$.
We are now in the position of proving Theorem 1.1.
Proof of Theorem 1.1.
Pick any $\mathbf{a}=(a_{n})_{n\geq 1}\in\overline{\mathsf{R}}$ and let $(\mathbf{b}^{n})_{n\geq 1}$ be a sequence in $\mathsf{R}$ converging to $\mathbf{a}$. Then, $\overline{\Lambda}(\mathbf{a})\in\overline{\mathfrak{F}}$ because $\overline{\Lambda}$ is continuous and $\overline{\Lambda}(\mathbf{b}^{n})=\Lambda(\mathbf{a})\in\mathfrak{F}$ for all $n\in\mathbb{N}$. Next, we show $\overline{\Lambda}(\mathbf{a})\in\overline{\mathfrak{F}}\setminus\in\mathbb{Q}(i)$ by obtaining infinitely many coprime pairs $(p,q)\in\mathbb{Z}[i]$ satisfying
(10) $\left|\overline{\Lambda}(\mathbf{a})-\frac{p}{q}\right|\leq\frac{1}{|q|^{2}}.$
Pick $N\in\mathbb{N}$ and write $p_{N}=p_{N}(a_{1},\ldots,a_{N})$, $q_{N}=q_{N}(a_{1},\ldots,a_{N})$. For all large $n\in\mathbb{N}$, we have $\operatorname{pref}(\mathbf{b}^{n};N)=\operatorname{pref}(\mathbf{a};N)$ and, by Proposition 3.1.v,
$\left|\overline{\Lambda}(\mathbf{a})-\frac{p_{N}}{q_{N}}\right|\leq\left|\overline{\Lambda}(\mathbf{a})-\overline{\Lambda}(\mathbf{b}^{n})\right|+\left|\overline{\Lambda}(\mathbf{b}^{n})-\frac{p_{N}}{q_{N}}\right|\leq\left|\overline{\Lambda}(\mathbf{a})-\overline{\Lambda}(\mathbf{b}^{n})\right|+\frac{1}{|q_{N}|^{2}}.$
Letting $n\to\infty$ and using the continuity of $\overline{\Lambda}$, we see that $(p,q)=(p_{N},q_{N})$ solves (10). Furthermore, by Proposition 3.1.ii and (6), the pairs $(p_{N},q_{N})$, $N\in\mathbb{N}$, are coprime and different. Therefore, $\overline{\Lambda}(\mathbf{a})\in\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)$.
Now take any $z\in\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)$. Assume that $z=[0;a_{1},a_{2},\ldots]\in\mathfrak{F}$. If $\mathbf{a}\in\mathsf{R}$, then $z=\overline{\Lambda}(\mathbf{a})$. When $\mathbf{a}\in\mathsf{Ir}$, the sequence $\mathbf{b}$ from Lemma 4.1 verifies $\Lambda(\mathbf{a})=\overline{\Lambda}(\mathbf{b})$. When $z\not\in\mathfrak{F}$, we have either $\operatorname{Re}(z)=\tfrac{1}{2}$ or $\operatorname{Im}(z)=\tfrac{1}{2}$. If $\operatorname{Im}(z)=\frac{1}{2}$, then $\overline{z}=\mathsf{mir}_{1}(z)\in\mathfrak{F}$ and there is some $\mathbf{b}=(b_{n})_{n\geq 1}\in\overline{\mathsf{R}}$ such that $\overline{z}=\Lambda(\mathbf{b})$. Thus, by Proposition 3.4, we have $\mathbf{c}:=(\mathsf{mir}_{1}(b_{j}))_{j\geq 1}\in\overline{\mathsf{R}}$ and $\overline{\Lambda}(\mathbf{c})=z$. A similar argument replacing $\mathsf{mir}_{1}$ with $\mathsf{mir}_{2}$ works when $\operatorname{Re}(z)=\frac{1}{2}$. ∎
Remark 4.2.
Theorem 5.3 in [9] implies that $\overline{\Lambda}:\overline{\mathsf{R}}\to\overline{\mathfrak{F}}$ is at most six-to-one. However, in Section 5 we describe a large set where it is one-to-one.
4.2. The algorithm
For each $a\in\mathscr{D}$ such that
$a\in\{1+im,-1+im,m+i,m-i\}\text{ for some }m\in\mathbb{Z},\;|m|\geq 2,$
define
$S(a)\colon=\begin{cases}im,\text{ if }a\in\{1+im,-1+im\},\\ m,\text{ if }a\in\{m+i,m-i\}.\end{cases}$
Algorithm 4.3.
  • Input. A sequence $\mathbf{a}=(a_{n})_{n\geq 1}\in\mathsf{Ir}$.
  • Output. A sequence $\mathbf{b}=(b_{j})_{j\geq 1}\in\overline{\mathsf{R}}$ such that $\overline{\Lambda}(\mathbf{b})=\Lambda(\mathbf{a})$.
  1. Step 1.
    Put $N=0$ and define $\mathbf{b}_{N}\colon=(b_{N}(j))_{j\geq 1}$ by
    $\forall j\in\mathbb{N}\quad b_{N}(j):=a_{j}.$
  2. Step 2.
    If possible, pick $j_{N}\in\mathbb{N}$ such that
    (11) $(b_{N}(1),\ldots,b_{N}(j_{N}))\in\mathsf{R}(j_{N})\quad\text{ and }\quad(b_{N}(1),\ldots,b_{N}(j_{N}+1))\not\in\mathsf{R}(j_{N}+1).$
    Define the sequence $\mathbf{b}_{N+1}=(b_{N+1}(j))_{j\geq 1}$ as follows:
    1. i.
      For $l\in\{1,\ldots,j_{N}\}$, write
      $b_{N+1}(l):=b_{N}(l).$
    2. ii.
      For $l=j_{N}+1$, write
      $b_{N+1}(j_{N}+1):=S\left(b_{N}(j_{N}+1)\right).$
    3. iii.
      For $l\in\{j_{N}+1+n:n\in\mathbb{N}\}$, let $m\in\mathbb{Z}$ be such that $b_{N}(j_{N}+1)\in\{1+im,-1+im,m+i,m-i\}$.
      • If $b_{N}(j_{N}+1)\in\{m+i,m-i\}$ for some $m\in\mathbb{Z}$, put
        $\forall n\in\mathbb{N}\quad b_{N+1}(j_{N}+1+n):=\mathsf{mir}_{1}\left(b_{N}(j_{N}+1+n)\right).$
      • If $b_{N}(j_{N}+1)\in\{1+im,-1+im\}$, then
        $\forall n\in\mathbb{N}\quad b_{N+1}(j_{N}+1+n):=\mathsf{mir}_{2}\left(b_{N}(j_{N}+1+n)\right).$
    If there is no $j_{N}\in\mathbb{N}$ such that (11) holds, define $\mathbf{b}_{N+1}\colon=\mathbf{b}_{N}$. Put $N:=N+1$.
  3. Step 3.
    Repeat Step 2.
  4. Step 4.
    Take $\mathbf{b}:=\displaystyle\lim_{N\to\infty}\mathbf{b}_{N}$. $\blacksquare$
4.3. Examples
For the sake of clarity, we illustrate the algorithm with a couple of examples.
Example 4.1.
Let $(B_{n})_{n\geq 1}$ be a sequence in $\mathbb{Z}$ such that $|B_{n}|\geq 3$ for all $n\in\mathbb{N}$ and let $\mathbf{a}=(a_{n})_{n\geq 1}$ be given by
$\mathbf{a}=(-2,1+iB_{1},-2,1+iB_{2},-2,1+iB_{3},-2,1+iB_{4},\ldots).$
First, let us show that $\mathbf{a}\in\mathsf{Ir}$. Since $|B_{1}|\geq 3$, we have
$\mathfrak{F}_{2}(-2,1+iB_{1})=\left[-\frac{1}{2}-\frac{i}{2},-\frac{1}{2}+\frac{i}{2}\right),$
so $(-2,1+iB_{1})\in\mathsf{Ir}(2)$ and $[\zeta_{2},\zeta_{1}]=\overline{\mathcal{C}}_{1}(-2)\cap\overline{\mathfrak{F}}_{2}(-2,1+iB_{1})\subseteq\mathfrak{F}_{2}(-2,1+iB_{1})$ (cfr. Figure 2), which implies
$\overline{\mathcal{C}}_{3}(-2,1+iB_{1},-2)\subseteq\mathcal{C}_{2}(-2,1+iB_{1}).$
Moreover, from $\mathfrak{F}_{3}(-2,1+iB_{1},-2)=\mathfrak{F}\cap C(1)$, we conclude
$\iota\left[\overline{\mathfrak{F}}_{3}(-2,1+iB_{1},-2)\right]\cap\tau_{1+iB_{2}}[\overline{\mathfrak{F}}]$ $=\left[1+i\left(B_{2}-\frac{1}{2}\right),1+i\left(B_{2}+\frac{1}{2}\right)\right]$
$\subseteq\iota\left[\mathfrak{F}_{3}(-2,1+iB_{1},-2)\right].$
and hence, by (8),
$\overline{\mathcal{C}}_{4}(-2,1+iB_{1},-2,1+iB_{2})\subseteq\mathcal{C}_{3}(-2,1+iB_{1},-2).$
We may continue this way to show
$\forall n\in\mathbb{N}\quad\overline{\mathcal{C}}_{n+1}(a_{1},\ldots,a_{n+1})\subseteq\mathcal{C}_{n}(a_{1},\ldots,a_{n}).$
Since $|\mathcal{C}_{n}(a_{1},\ldots,a_{n})|\to 0$ as $n\to\infty$, we can single out a complex number $z$ satisfying
$\{z\}=\bigcap_{n\in\mathbb{N}}\mathcal{C}_{n}(a_{1},\ldots,a_{n}).$
As a consequence, $z=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}$ and $\mathbf{a}\in\mathsf{Ir}$, because $(-2,1+iB_{1})\in\mathsf{Ir}(2)$. The algorithm applied to $\mathbf{a}$ gives us the sequence
$\mathbf{b}=(b_{j})_{j\geq 1}=(-2,iB_{1},2,iB_{2},-2,iB_{3},2,iB_{4},-2,iB_{5},2,iB_{6},\ldots),$
To see it, we work out the first two iterations.
  1. i.
    Take $N=0$ and $\mathbf{b}_{0}\colon=\mathbf{a}$. Then, $j_{0}=1$ and
    $b_{1}(1):=a_{1}=-2,\quad b_{1}(2):=S(a_{2})=iB_{1},\quad\forall l\in\mathbb{N}_{\geq 3}\;\,b_{1}(l):=\mathsf{mir}_{2}(a_{l}),$
    which means
    $\mathbf{b}_{1}=(-2,iB_{1},2,-1+iB_{2},2,-1+iB_{3},2,-1+iB_{4},\ldots).$
    Add $1$ to $N$.
  2. ii.
    For $N=1$, we have $j_{1}=3$ because
    $(-2,iB_{1},2)\in\mathsf{R}(3)\quad\text{ and }\quad(-2,iB_{1},2,-1+iB_{2})\not\in\mathsf{R}(4).$
    Note that $(-2,iB_{1},2,-1+iB_{2})\not\in\Omega(4)$. We define
    $\forall l\in\{1,2,3\}\quad b_{2}(l)$ $:=b_{1}(l),$
    $b_{2}(4)$ $:=S(b_{1}(3))=iB_{2},$
    $\forall l\in\mathbb{N}_{\geq 5}\quad b_{2}(l)$ $:=\mathsf{mir}_{2}(b_{1}(l))=\mathsf{mir}_{2}^{2}(a_{l})=a_{l},$
    which implies
    $\mathbf{b}_{2}=(-2,iB_{1},2,iB_{2},-2,1+iB_{3},-2,1+iB_{4},\ldots).$
We may replace the condition $\min\{|B_{n}|:n\in\mathbb{N}\}=3$ by the optimal condition $\min\{|B_{n}|:n\in\mathbb{N}\}=2$ at the expense of some notational complications.
Example 4.2.
Algorithm 4.3 applied on the extremely irregular sequence
$\mathbf{a}=(-2,1+2i,-2+i,1+2i,-2+i,1+2i,-2+i,\ldots),$
where the word $(1+2i,-2+i,1+2i,-2+i)$ is repeated, gives
$\mathbf{b}=(-2,2i,2,-2i,-2,2i,2,-2i,\ldots).$
Observe that $\zeta_{1}=\overline{\Lambda}(\mathbf{a})$ and that the sequences
[…]
also belong to $\overline{\mathsf{R}}$ and are mapped to $\zeta_{1}$ by $\overline{\Lambda}$ (cfr. [9, Proposition 5.5]).
5. First application: Borel hierarchy of the set of Hurwitz normal numbers
The objective of this section is to analyse the complexity of the set of Hurwitz normal numbers with respect to the Borel hierarchy employed in descriptive set theory. This hierarchy provides insight into the complexity for determining whether a complex number is a Hurwitz normal number. The main step of this proof is to obtain dynamical information of the subshift $\overline{\mathsf{R}}$ (Theorem 1.3). Subsequently, we leverage a result from Airey, Jackson, Kwietniak, and Mance to establish Theorem 5.1 below, which implies Theorem 1.2.
5.1. Hurwitz normal numbers
In [31], Nakada proved the existence and uniqueness of a $T$-ergodic Borel probability measure $\mu_{\operatorname{H}}$ on $\mathfrak{F}$ that is equivalent to the Lebesgue measure. By virtue of Birkhoff’s Ergodic Theorem [37, Theorem 1.14], for every $n\in\mathbb{N}$ and every $\mathbf{b}\in\mathsf{R}(n)$, almost every $z\in\mathfrak{F}$ with respect to $\mu_{\operatorname{H}}$ (or, equivalently, to Lebesgue measure) satisfies
(12) $\lim_{n\to\infty}\frac{\#\{j\in\{0,\ldots,n-1\}:(a_{j}(z),\ldots,a_{j+n-1}(z))=\mathbf{b}\}}{n}=\mu_{\operatorname{H}}(\mathcal{C}_{n}(\mathbf{b})).$
In fact, for almost every $z\in\mathfrak{F}$ the equality (12) holds for every finite regular word $\mathbf{b}$. We refer to such numbers $z$ as Hurwitz normal numbers and we denote them by $\operatorname{Norm}(\mu_{\operatorname{H}})$.
As for regular continued fractions or integer base expansions, normality is closely related to equidistribution. Given a sequence $(z_{n})_{n\geq 1}$ in $\mathfrak{F}$, let $\delta_{z_{n}}$ be the point mass measure on $z_{n}$ for each $n\in\mathbb{N}$. We call $(z_{n})_{n\geq 1}$ uniformly distributed$\pmod{\mu_{\operatorname{H}}}$ if the sequence of measures $\frac{1}{n}\sum_{j=1}^{n}\delta_{z_{n}}$ converges weakly to $\mu_{\operatorname{H}}$. From the Portmanteau Theorem [23, Theorem 5.25], we can derive the next equality:
$\operatorname{Norm}(\mu_{\operatorname{H}})=\left\{z\in\mathfrak{F}:(T^{n}(z))_{n\geq 1}\text{ is uniformly distributed}\pmod{\mu_{\operatorname{H}}}\right\}.$
In particular, for every $z\in\operatorname{Norm}(\mu_{\operatorname{H}})$ the set $\{T^{n}(z):n\in\mathbb{N}_{0}\}$ is dense in $\mathfrak{F}$.
5.2. Borel hierarchy
The statements in this subsection are classical results in descriptive set theory, their proofs can be found in [26, Chapter 11]. Let $(X,\eta)$ be a Polish space; that is, a separable completely metrizable topological space. Denote the first uncountable ordinal by $\omega_{1}$. We define $\Sigma_{1}^{0}(X):=\eta$ and for each ordinal $\xi$ with $1\leq\xi<\omega_{1}$ put
$\mathbf{\Pi}_{\xi}^{0}(X)$ :=$\colon=\left\{X\setminus A\colon A\in\mathbf{\Sigma}_{\xi}^{0}(X)\right\}\text{, and}$
$\mathbf{\Sigma}_{\xi}^{0}(X)$ :=$\colon=\left\{\bigcup_{n\in\mathbb{N}}A_{n}:\forall n\in\mathbb{N}\quad\exists\xi_{n}<\xi\;\left(A_{n}\in\mathbf{\Pi}_{\xi_{n}}^{0}(X)\right)\right\}.$
In other words, $\Sigma_{1}^{0}(X)$ is the collection of open sets, $\Pi_{1}^{0}(X)$ the collection of closed sets, $\Sigma_{1}^{0}(X)$ the collection of countable unions of closed sets (an example is $\mathbb{Q}$ when $X=\mathbb{R}$), and so on. This family of collections has a hierarchical structure; that is, for any two ordinals $\xi,\xi^{\prime}$ satisfying $1\leq\xi<\xi^{\prime}<\omega_{1}$ we have
$\mathbf{\Pi}_{\xi}^{0}(X)\cup\mathbf{\Sigma}_{\xi}^{0}(X)\subseteq\mathbf{\Pi}_{\xi^{\prime}}^{0}(X)\quad\text{ and }\quad\mathbf{\Pi}_{\xi}^{0}(X)\cup\mathbf{\Sigma}_{\xi}^{0}(X)\subseteq\mathbf{\Sigma}_{\xi^{\prime}}^{0}(X).$
Furthermore, every Borel subset belongs to the family, that is,
$\mathscr{B}(X)=\bigcup_{\xi<\omega_{1}}\mathbf{\Pi}_{\xi}^{0}(X)=\bigcup_{\xi<\omega_{1}}\mathbf{\Sigma}_{\xi}^{0}(X),$
where $\mathscr{B}(X)$ is the Borel $\sigma$-algebra of $(X,\eta)$. This family of collections is known as the Borel hierarchy.
Note that whenever $X$ is uncountable, all the collections $\mathbf{\Pi}_{\xi}^{0}(X)$ and $\mathbf{\Sigma}_{\xi}^{0}(X)$ are different. The Borel hierarchy is a standard measurement of the complexity of Borel subsets, and it is related to the number of quantifiers needed to describe a subset. For this reason, it is important to know which is the first collection where a Borel set appears. For any ordinal $\xi<\omega_{1}$, we say that a set $A\in\mathscr{B}(X)$ is $\mathbf{\Pi}_{\xi}^{0}(X)$-hard if $A\not\in\mathbf{\Sigma}_{\xi}^{0}(X)$ We say that $A\in\mathscr{B}(X)$ is $\mathbf{\Pi}_{\xi}^{0}(X)$-complete if it belongs to $\mathbf{\Pi}_{\xi}^{0}(X)$ and it is $\mathbf{\Pi}_{\xi}^{0}(X)$-hard. Note that if $A$ is $\mathbf{\Pi}_{\xi}^{0}(X)$-complete, then $A\in\mathbf{\Pi}_{\xi^{\prime}}^{0}(X)\cap\mathbf{\Sigma}_{\xi^{\prime}}^{0}(X)$ for every countable ordinal $\xi^{\prime}>\xi$.
We are ready to state the main result of this section.
Theorem 5.1.
The set $\operatorname{Norm}(\mu_{\operatorname{H}})$ is $\mathbf{\Pi}_{3}^{0}(\mathbb{C})$-complete.
We will use a variant of the specification to prove this result.
5.3. Feeble specification
Let $\mathscr{A}$ be non-empty and at most countable set. We denote non-empty words $\mathbf{u}$ on $\mathscr{A}$ as $\mathbf{u}=(u_{1},u_{2},\ldots)$. In symbolic dynamics, it is customary to represent words without parentheses, i.e., $\mathbf{u}=u_{1}u_{2}\cdots$, however, we decided to keep them.
For any two non-empty finite words $\mathbf{v}=(v_{1},\ldots,v_{n})$, $\mathbf{w}=(w_{1},\ldots,w_{n})$ of the same length $n\in\mathbb{N}$, the normalized Hamming distance between $\mathbf{v}$ and $\mathbf{w}$ is
$d_{H}(\mathbf{v},\mathbf{w})=\frac{\#\{j\in\{1,\ldots,n\}:v_{j}\neq w_{j}\}}{n}.$
Endow $\mathscr{A}^{\mathbb{N}}$ with the product topology. A subshift $X\subset\mathscr{A}^{\mathbb{N}}$ is a closed subset such that $\sigma[X]=X$. Given a subshift $X$, we denote by $\mathscr{L}(X)$ the set of all finite words that appear as factors of points in $X$. Given $\mathbf{b}\in\mathscr{L}(\mathscr{A}^{\mathbb{N}})\setminus\{\epsilon\}$, we define the symbolic cylinder based on $\mathbf{b}$ as
$[\mathbf{b}]_{X}\colon=\left\{\mathbf{a}\in X:\operatorname{pref}(\mathbf{a};|\mathbf{b}|)=\mathbf{b}\right\}.$
The subshift $X$ has the right feeble specification property if there is a set $\mathscr{G}\subseteq\mathscr{L}(X)$ with the following properties.
  1. i.
    If $\mathbf{u},\mathbf{v}\in\mathscr{G}$, then $\mathbf{u}\mathbf{v}\in\mathscr{G}$.
  2. ii.
    For every $\varepsilon>0$ there exists $N\in\mathbb{N}$ such that for any $\mathbf{u}\in\mathscr{G}$ and any $\mathbf{v}\in\mathscr{L}(X)$ satisfying $|\mathbf{v}|\geq N$ there are some $\mathbf{s}^{\prime},\mathbf{v}^{\prime}\in\mathscr{L}(\mathscr{A}^{\mathbb{N}})$ satisfying
    $|\mathbf{v}^{\prime}|=|\mathbf{v}|,\quad 0\leq|\mathbf{s}|\leq\varepsilon|\mathbf{v}|,\quad d_{H}(\mathbf{v},\mathbf{v}^{\prime})<\varepsilon,\quad\mathbf{u}\mathbf{s}\mathbf{v}^{\prime}\in\mathscr{G}.$
For any $\mathbf{x}=(x_{1},x_{2},x_{3},\ldots)\in\mathscr{A}^{\mathbb{N}}$ and $\mathbf{w}\in\mathscr{L}(\mathscr{A}^{\mathbb{N}})$ define
$e(\mathbf{w},\mathbf{x},N)\colon=\#\left\{j\in\{1,\ldots,N\}:(x_{j},x_{j+1},\ldots,x_{j+n-1})=\mathbf{w}\right\}.$
The quasi-regular set $Q(X)$ of $X$ is the set of words for which the frequency of every finite word exists:
$Q(X)\>=\left\{\mathbf{x}\in X:\forall\mathbf{w}\in\mathscr{L}(X)\quad\lim_{N\to\infty}\frac{e(\mathbf{w},\mathbf{x},N)}{N}\text{ exists }\right\}.$
For a $\sigma$-invariant probability measure $\nu$, we say that a word $\mathbf{x}\in X$ is $\nu$-generic if
$\forall\mathbf{w}\in\mathscr{L}(\mathscr{A}^{\mathbb{N}})\quad\lim_{N\to\infty}\frac{e(\mathbf{w},\mathbf{x},N)}{N}=\nu\left([\mathbf{w}]_{X}\right).$
We denote by $G_{\nu}$ the set of $\nu$-generic points. Clearly, $Q(X)\subseteq G_{\nu}$. By Birkhoff’s ergodic theorem, if $\nu$ is an $\sigma$-ergodic probability measure, then almost every point is $\nu$-generic.
In [4], Airey, Jackson, Kwietniak, and Mance developed a tool to determine the level in the Borel hierarchy of the set of generic sets of subshifts with the feeble specification. They applied this result to study normal numbers of regular continued fractions and other numeration systems.
Theorem 5.2 ([4, Theorem 6]).
Let $X$ be a subshift with the right feeble specification property and at least two $\sigma$-invariant Borel probability measures. Let $\nu$ be a shift-invariant Borel probability measure. If $B$ is a Borel set such that $G_{\nu}\subseteq B\subseteq Q(X)$, then $B$ is $\mathbf{\Pi}_{3}^{0}(X)$-hard. In particular, $G_{\nu}$ and $Q(X)$ are $\mathbf{\Pi}_{3}^{0}(X)$-complete.
5.4. Proof of Theorem 1.3
The next lemma follows from [13, Proposition 8.4].
Lemma 5.3.
Let $n\in\mathbb{N}$ and $\mathbf{a}=(a_{1},\ldots,a_{n})\in\mathsf{R}(n)$ be arbitrary.
  1. i.
    If $\operatorname{Pm}(a_{n})\geq 3$ and $\mathbf{b}\in\mathsf{R}(m)$ for some $m\in\mathbb{N}$, then $\mathbf{a}\mathbf{b}\in\mathsf{R}(m+n)$.
  2. ii.
    There exists some $b\in\mathbb{Z}[i]$ such that $\operatorname{Pm}(b)\geq 3$ and $\mathfrak{F}_{n+1}(\mathbf{a}b)=\mathfrak{F}$.
Proof of Theorem 1.3.
By definition, $\overline{\mathsf{R}}$ is closed in $\mathscr{D}^{\mathbb{N}}$. Since factors (subwords) of regular words are also regular, $\sigma[\mathsf{R}]\subseteq\mathsf{R}$. If $\mathbf{a}\in\mathsf{R}$ and $b=3+3i$, then $b\mathbf{a}\in\mathsf{R}$ and $\sigma(b\mathbf{a})=\mathbf{a}$, so $\mathsf{R}\subseteq\sigma[\mathsf{R}]$. Define the set
$\mathscr{G}\colon=\left\{(c_{1},\ldots,c_{m})\in\bigcup_{n\in\mathbb{N}}\mathsf{R}(n):\operatorname{Pm}(c_{m})\geq 3\right\}.$
Let us verify the conditions defining the right feeble specification property.
  1. i.
    If $\mathbf{u},\mathbf{v}\in\mathscr{G}$, then $\mathbf{u}\mathbf{v}\in\mathscr{G}$ by Lemma 5.3.i.
  2. ii.
    Let $\varepsilon>0$ and $N\in\mathbb{N}$ such that $1/N<\varepsilon$. Take $\mathbf{v}=(v_{1},\ldots,v_{m})\in\mathscr{L}(\overline{\mathsf{R}})$ with $m\geq N$. Then $\mathbf{v}\in\mathsf{R}(m)$, so $(v_{1},\ldots,v_{m-1})\in\mathsf{R}(m-1)$ and, by Lemma 5.3.ii, there exists $a\in\mathscr{D}$ satisfying
    $\operatorname{Pm}(a)\geq 3\quad\text{ and }\quad\mathbf{v}^{\prime}\colon=(v_{1},\ldots,v_{m-1},a)\in\mathscr{G}.$
    Letting $\mathbf{s}$ be the empty word, for every $\mathbf{u}\in\mathscr{G}$ we have
    $|\mathbf{s}|\leq\varepsilon|\mathbf{v}|,\quad|\mathbf{v}^{\prime}|=|\mathbf{v}|,\quad d_{H}(\mathbf{v},\mathbf{v}^{\prime})<\varepsilon\quad\mathbf{u}\mathbf{s}\mathbf{v}^{\prime}\in\mathscr{G}.\;\qed$
5.5. Proof of Theorem 5.1
We split the proof of Theorem 5.1 into Lemma 5.5 (lower bound) and Lemma 5.6 (upper bound).
Let $X,Y$ be Polish spaces. We call $A\subseteq X$ Wadge reducible to $B\subseteq Y$ if there exists a continuous function $f:X\to Y$ (called a Wadge reduction) such that $f^{-1}[B]=A$. For more on Wadge reductions, see [26, Section 21.E].
Remark 5.4.
Using basic properties of continuous functions, one can show that if a set $A$ is $\mathbf{\Pi}_{3}^{0}(X)$-hard and $A\subset X$ is Wadge reducible to $B\subset Y$, then $B$ is $\mathbf{\Pi}_{3}^{0}(Y)$-hard.
Lemma 5.5.
The set $\operatorname{Norm}(\mu_{\operatorname{H}})$ is $\mathbf{\Pi}_{3}^{0}(\mathbb{C})$-hard.
Proof.
By Proposition 3.3, any constant sequence $\mathbf{a}=(a,a,\ldots)\in\mathscr{D}^{\mathbb{N}}$ with $|a|\geq\sqrt{8}$ belongs to $\overline{\mathsf{R}}$ and the point mass measure based on $\mathbf{a}$ is $\sigma$-invariant. This implies that $(\overline{\mathsf{R}},\sigma)$ has infinitely many $\sigma$-invariant measures.
Let us construct the $\sigma$-invariant Borel probability measure $\nu$ on $\overline{\mathsf{R}}$ on which we apply Theorem 5.2. We define the set
$\mathscr{I}\colon=\bigcap_{n\in\mathbb{N}}\bigcup_{\mathbf{a}\in\mathsf{R}(n)}\mathcal{C}^{\circ}_{n}(\mathbf{a})\subseteq\mathfrak{F}.$
It is clear that $\mathscr{I}\subset\mathbb{C}$ is a Borel subset. We will use the functions $\Lambda$ and $\overline{\Lambda}\colon\overline{\mathsf{R}}\to\mathbb{C}$ defined in Subsection 4.1. By definition $\overline{\Lambda}$ is continuous.
We define the Borel probability measure $\nu$ on $\overline{\mathsf{R}}$ as follows.
  1. i.
    We set $\nu(B):=\mu_{\operatorname{H}}(\overline{\Lambda}[B])$ for every Borel subset $B\subset\overline{\Lambda}\,^{-1}[\mathscr{I}].$
  2. ii.
    $\nu\left(\overline{\mathsf{R}}\setminus\overline{\Lambda}\,^{-1}[\mathscr{I}]\right)=0$.
Since $\mu_{\operatorname{H}}$ is $T$-invariant and $\overline{\Lambda}\circ\sigma=T\circ\overline{\Lambda}$ (Equation (4)), we have that $\nu$ is $\sigma$-invariant. Theorem 5.2 implies that $G_{\nu}\subseteq\overline{\mathsf{R}}$ is $\Pi_{3}^{0}(\overline{\mathsf{R}})$-complete.
To finish the proof, we must translate this conclusion to the complex plane using $\overline{\Lambda}$ as a Wadge reduction. First, note that if $z$ belongs to the boundary of some prototype set, then $T(z)$ also belongs to the boundary of some prototype set (the proof of this assertion amounts to applying the complex inversion to a finite collection of line segments and arches). Since for each $z\in\operatorname{Norm}(\mu_{\operatorname{H}})$ the set $\{T^{n}(z):n\in\mathbb{N}_{\geq 0}\}$ is dense in $\mathfrak{F}$, we have $\operatorname{Norm}(\mu_{\operatorname{H}})\subseteq\mathscr{I}$. Hence, the definitions of $G_{\nu}$ and $\operatorname{Norm}(\mu_{\operatorname{H}})$ yield
$G_{\nu}=\overline{\Lambda}\,^{-1}[\operatorname{Norm}(\mu_{\operatorname{H}})].$
Therefore $G_{\nu}\subseteq\overline{\mathsf{R}}$ is Wadge reducible to $\operatorname{Norm}(\mu_{\operatorname{H}})\subseteq\mathbb{C}$. Using Remark 5.4 we conclude that $\operatorname{Norm}(\mu_{\operatorname{H}})$ is $\mathbf{\Pi}_{3}^{0}(\mathbb{C})$-hard. ∎
Lemma 5.6.
The set $\operatorname{Norm}(\mu_{\operatorname{H}})$ is $\mathbf{\Pi}_{3}^{0}(\mathbb{C})$.
Proof.
Let $(\varepsilon_{k})_{k\geq 1}$ be a strictly decreasing sequence of positive numbers converging to 00. Given $A\subseteq\mathfrak{F}$, let $\chi_{A}$ be its indicator function. For $M,n,k\in\mathbb{N}$ and $\mathbf{b}\in\mathsf{R}(n)$, define the sets
$E_{n}(\mathbf{b},M,k)$ :=$:=\left\{z=[0;a_{1},a_{2},\ldots]:\left|\frac{1}{M}\sum_{j=0}^{M-1}\chi_{\mathcal{C}_{n}(\mathbf{b})}\left(T^{j}(z)\right)-\mu_{\operatorname{H}}(\mathcal{C}_{n}(\mathbf{a}))\right|\leq\varepsilon_{k}\right\},$
$F_{n}(\mathbf{b},M,k)$ $:=\operatorname{Cl}_{\mathbb{C}}(E_{n}(\mathbf{b},M,k))\text{, and}$
$G_{n}(\mathbf{b},M,k)$ $:=\operatorname{Int}_{\mathbb{C}}(E_{n}(\mathbf{b},M,k)).$
Note that each $E_{n}(\mathbf{b},M,k)$ is a finite union of cylinders of level $n+M-1$. The definition of $\operatorname{Norm}(\mu_{\operatorname{H}})$ gives
$\operatorname{Norm}(\mu_{\operatorname{H}})=\bigcap_{n\in\mathbb{N}}\bigcap_{\mathbf{b}\in\mathsf{R}(n)}\bigcap_{k\in\mathbb{N}}\bigcup_{N\in\mathbb{N}}\bigcap_{M\in\mathbb{N}_{\geq N}}E_{n}(\mathbf{b},M,k).\$
Put
$\mathfrak{L}$ $\colon=\bigcap_{n\in\mathbb{N}}\bigcap_{\mathbf{b}\in\mathsf{R}(n)}\bigcap_{k\in\mathbb{N}}\bigcup_{N\in\mathbb{N}}\bigcap_{M\in\mathbb{N}_{\geq N}}G_{n}(\mathbf{b},M,k),$
$\mathfrak{M}$ $\colon=\bigcap_{n\in\mathbb{N}}\bigcap_{\mathbf{b}\in\mathsf{R}(n)}\bigcap_{k\in\mathbb{N}}\bigcup_{N\in\mathbb{N}}\bigcap_{M\in\mathbb{N}_{\geq N}}F_{n}(\mathbf{b},M,k),$
so $\mathfrak{L}\subseteq\operatorname{Norm}(\mu_{\operatorname{H}})\subseteq\mathfrak{M}$. Note that $\bigcup_{N\in\mathbb{N}}\bigcap_{M\in\mathbb{N}_{\geq N}}F_{n}(\mathbf{b},M,k)\in\mathbf{\Sigma}_{2}^{0}(\mathbb{C})$, this implies that $\mathfrak{M}\in\mathbf{\Pi}_{3}^{0}(\mathbb{C})$. We will show that $\mathfrak{M}\subseteq\mathfrak{L}$ to conclude the lemma. Let $z\in\mathfrak{M}$ be arbitrary and assume that
(13) $z\in\bigcap_{n\in\mathbb{N}}\mathcal{C}^{\circ}_{n}\left(a_{1}(z),\ldots,a_{n}(z)\right).$
Let $n,k\in\mathbb{N}$, $\mathbf{b}\in\mathsf{R}(n)$ and consider $N\in\mathbb{N}$ with
$\forall M\in\mathbb{N}_{\geq N}\quad z\in F_{n}(\mathbf{b},M,k).$
Take any $M\in\mathbb{N}_{\geq N}$ and let $\mathbf{d}\in\mathsf{R}(M+n)$ be such that
$\mathcal{C}_{M+n}(\mathbf{d})\subseteq E_{n}(\mathbf{b},M,k)\quad\text{ and }\quad z\in\overline{\mathcal{C}}_{M+n}(\mathbf{d}).$
Since different regular cylinders of the same level have disjoint interior, from (13) we get
$\mathbf{d}=\left(a_{1}(z),\ldots,a_{M+n}(z)\right),$
so $z\in G_{n}(\mathbf{b},M,k)$. We conclude $\mathfrak{M}\subseteq\mathfrak{L}$, hence $\operatorname{Norm}(\mu_{\operatorname{H}})=\mathfrak{M}$, if (13) holds for all $z\in\mathfrak{M}$.
Lastly, we show that (13) is indeed true for all $z\in\mathfrak{M}$. To this end, pick any $z\in\mathfrak{M}$. It suffices to prove that $\{T^{n}(z):n\in\mathbb{N}_{\geq 0}\}$ is dense in $\mathfrak{F}$, which in turns follows from
(14) $\forall n\in\mathbb{N}\quad\forall\mathbf{b}\in\mathsf{R}(n)\quad\{T^{n}(z):n\in\mathbb{N}_{\geq 0}\}\cap\mathcal{C}^{\circ}_{n}(\mathbf{b})\neq\varnothing.$
Take any $n\in\mathbb{N}$ and $\mathbf{b}\in\mathsf{R}(n)$. Let $d\in\mathscr{D}$ be such $\operatorname{Pm}(d)\geq 3$ and $c\in\mathscr{D}$ such that
$\mathbf{b}c\in\mathsf{R}(n+1),\quad\text{ and }\quad\overline{\mathcal{C}}_{n+1}(\mathbf{b}c)\subseteq\mathcal{C}^{\circ}_{n}(\mathbf{b})$
(see Lemma 7.1 below). Since $d\mathbf{b}c\in\mathsf{R}(n+2)$ by Lemma 5.3. i, we may choose $k\in\mathbb{N}$ so large that
$\varepsilon_{k}<\frac{\mu_{\operatorname{H}}(\mathcal{C}_{n+1}(d\mathbf{b}c))}{4}$
and $N\in\mathbb{N}$ with
$\forall M\in\mathbb{N}_{\geq N}\quad z\in F_{n+2}(d\mathbf{b}c,M,k).$
Take $\mathbf{d}\in\mathsf{R}(M+n-1)$ satisfying
$\mathcal{C}_{M+n-1}(\mathbf{d})\subseteq E_{n+2}(d\mathbf{b}c,M,k)\quad\text{ and }\quad z\in\overline{\mathcal{C}}_{M+n-1}(\mathbf{d}).$
The choice of $k$ and the definition of $E_{n+1}(\mathbf{b}c,M,k)$ guaranty the existence of some $r\in\{n+1,\ldots,M+n\}$ satisfying
$(d_{r-n-1},\ldots,d_{r-1},d_{r})=(d,b_{1},\ldots,b_{n},c).$
Let us show that
(15) $\overline{\mathcal{C}}_{M+n-1}(\mathbf{d})\subseteq\overline{\mathcal{C}}_{r}(d_{1},\ldots,d_{r})\subseteq\mathcal{C}^{\circ}_{r-1}(d_{1},\ldots,d_{r-1}).$
The first contention is trivial. From
$\mathcal{C}_{r}(\mathbf{d})$ $=\iota\tau_{d_{1}}\cdots\iota\tau_{d_{r}}[\mathfrak{F}_{r}(d_{1},\ldots,d_{r})],$
$\mathcal{C}_{r-1}(d_{1},\ldots,d_{r-1})$ $=\iota\tau_{d_{1}}\cdots\iota\tau_{d_{r-1}}[\mathfrak{F}_{r-1}(d_{1},\ldots,d_{r-1})].$
and the fact that $\iota$ and the functions $\tau_{a}$, $a\in\mathscr{D}$, are homeomorphisms from the Riemann sphere onto itself, the second contention in (15) is equivalent to
(16) $\iota\tau_{c}\left[\overline{\mathfrak{F}}_{r}(\mathbf{d}^{\prime}d\mathbf{b}c)\right]\subseteq\mathfrak{F}^{\circ}_{r}(\mathbf{d}^{\prime}d\mathbf{b})\;\text{ for }\;\mathbf{d}^{\prime}=(d_{1},\ldots,d_{r-n-2}).$
Since $\operatorname{Pm}(d)\geq 3$, we may use Proposition 3.2 and induction on $|\mathbf{b}|$ that $\mathfrak{F}_{r}(\mathbf{d}^{\prime}d\mathbf{b}c)=\mathfrak{F}_{r}(\mathbf{b}c)$ and $\mathfrak{F}_{r-1}(\mathbf{d}^{\prime}d\mathbf{b})=\mathfrak{F}_{n}(\mathbf{b})$. Therefore, the second contention is (15) is equivalent to $\iota\tau_{c}[\overline{\mathfrak{F}}_{n+1}(\mathbf{b}c)]\subseteq\mathfrak{F}^{\circ}_{n}(\mathbf{b})$, which is equivalent to $\overline{\mathcal{C}}_{n+1}(\mathbf{b}c)\subseteq\mathcal{C}^{\circ}_{n}(\mathbf{b})$. We conclude $T^{r-n-1}(z)\in\mathcal{C}^{\circ}_{n}(\mathbf{b})$ and (14) follows. ∎
6. Second application: construction of transcendental numbers
Let $\mathscr{A}$ be a finite alphabet and $\mathbf{a}=(a_{n})_{n\geq 1}$ a sequence on $\mathscr{A}$. Recall that for each $n\in\mathbb{N}$ we have defined
$r(n,\mathbf{a}):=\min\left\{m\in\mathbb{N}:\exists j\in\{1,\ldots,m-n\}\quad(a_{j},\ldots,a_{j+n-1})=(a_{m},\ldots,a_{m+n-1})\right\}.$
The repetition exponent of $\mathbf{a}$ is
$\operatorname{rep}(\mathbf{a}):=\liminf_{n\to\infty}\frac{r(n,\mathbf{a})}{n}.$
Bugeaud and Kim [10] showed that the finiteness of $\operatorname{rep}(\mathbf{a})$ for a given sequence $\mathbf{a}$ is equivalent to a combinatorial condition used in [1, 8] to construct real transcendental numbers.
Lemma 6.1 ([10, Section 10]).
Let $\mathscr{A}$ be a finite alphabet and $\mathbf{a}=(a_{n})_{n\geq 1}\in\mathscr{A}^{\mathbb{N}}$. If $\operatorname{rep}(\mathbf{a})<\infty$, then there are three sequences $(W_{n})_{n\geq 1}$, $(U_{n})_{n\geq 1}$, $(V_{n})_{n\geq 1}$ of finite words in $\mathscr{A}$ such that
  1. 1.
    For all $n\in\mathbb{N}$, the word $W_{n}U_{n}V_{n}U_{n}$ is a prefix of $\mathbf{a}$;
  2. 2.
    The sequence $\left((|W_{n}|+|V_{n}|)/|U_{n}|\right)_{n\geq 1}$ is bounded above;
  3. 3.
    The sequence $(|U_{n}|)_{n\geq 1}$ is increasing.
Theorem 6.2 below is a complex version of Bugeaud’s transcendence criterion for non-periodic continued fractions with finite repetition exponent [8].
Theorem 6.2 ([19, Theorem 5.1]).
Let $\mathbf{a}\in\Omega$ be a bounded and non-periodic sequence such that $\operatorname{rep}(\mathbf{a})<\infty$. Let $(W_{n})_{n\geq 1}$, $(U_{n})_{n\geq 1}$, and $(V_{n})_{n\geq 1}$ be sequences as in Lemma 6.1.
  1. 1.
    If $\liminf_{n\to\infty}|W_{n}|<\infty$, then $\Lambda(\mathbf{a})$ is transcendental.
  2. 2.
    If $\liminf_{n\to\infty}|W_{n}|=\infty$ and $|a_{n}|\geq\sqrt{8}$ for all $n\in\mathbb{N}$, then $\Lambda(\mathbf{a})$ is transcendental.
We conjecture that the condition $|a_{n}|\geq\sqrt{8}$ for all $n\in\mathbb{N}$ can be removed in Theorem 6.2.2. The second application of our work, Theorem 1.4, is a construction of a family of transcendental numbers overlooked by Theorem 6.2. This provides evidence supporting our conjecture. A key ingredient of the proof is Schmidt’s extension of his celebrated Subspace Theorem to number fields, we only state it for $\mathbb{Q}(i)$. For $m\in\mathbb{N}$ and $\mathbf{z}=(z_{1},\ldots,z_{m})\in\mathbb{C}^{m}$, we write $\|\mathbf{z}\|=\max\{|z_{1}|,\ldots,|z_{m}|\}$ and $\overline{\mathbf{z}}=(\overline{z_{1}},\ldots,\overline{z_{m}})$.
Theorem 6.3 ([35, Theorem 3]).
Let $m\in\mathbb{N}$. Assume that we have two collections of linearly independent linear forms $\mathscr{L}_{1},\ldots,\mathscr{L}_{m}$ and $\mathscr{M}_{1},\ldots,\mathscr{M}_{m}$ in $m$ variables with real or complex algebraic coefficients. Let $\varepsilon>0$. Then, there are finitely many proper subspaces $T_{1},\ldots,T_{k}$ of $\mathbb{C}^{m}$ containing all the solutions $\mathbf{z}=(z_{1},\ldots,z_{m})\in\mathbb{Z}[i]^{m}$ of
$\left|\mathscr{L}_{1}(\mathbf{z})\cdots\mathscr{L}_{m}(\mathbf{z})\right|\left|\mathscr{M}_{1}(\overline{\mathbf{z}})\cdots\mathscr{M}_{m}(\overline{\mathbf{z}})\right|<\|\mathbf{z}\|^{-\varepsilon}.$
Recall that if $A_{0},A_{1},\ldots,A_{n}$ are variables and $q_{0},q_{1},\ldots,q_{n}$ are the denominators of the continued fraction $[A_{0};A_{1},A_{2},A_{3},\ldots,A_{n}]$, then
(17) $\frac{q_{n-1}}{q_{n}}=[0;A_{n},A_{n-1},\ldots,A_{2},A_{1}]$
(see, for example, [27, Theorem 6]). Given $\mathbf{B}\in\Omega$ as in Theorem 1.4, let $(W_{n})_{n\geq 1}$, $(U_{n})_{n\geq 1}$, $(V_{n})_{n\geq 1}$ be as in Lemma 6.1 and define the sequences of non-negative integers $(w_{n})_{n\geq 1}$, $(u_{n})_{n\geq 1}$, and $(v_{n})_{n\geq 1}$ by
$\forall n\in\mathbb{N}\quad w_{n}:=|W_{n}|,\;u_{n}:=|U_{n}|,\;v_{n}:=|V_{n}|,\;t_{n}:=w_{n}+u_{n}+v_{n}.$
The following lemma is essentially [19, Lemma 5.2].
Lemma 6.4.
Let $\zeta,\mathbf{B}$ be as in Theorem 1.4 and let $(q_{n})_{n\geq 1}$ correspond to $\zeta$ as in Equation (5). Then, there exists an $\varepsilon>0$ such that for all $n\in\mathbb{N}$ we have $\psi^{2u_{n}}\geq|q_{2w_{n}}{q_{2w_{n}+2u_{n}+2v_{n}}}|^{\varepsilon}$.
For any two finite words of the same length $A=(a_{1},\ldots,a_{n})$ and $B=(b_{1},\ldots,b_{n})$ the word $s(A,B):=(c_{1},\ldots,c_{2n})$ is given by $c_{2j-1}=a_{j}$ and $c_{2j}=b_{j}$ for all $j\in\{1,\ldots,n\}$; that is,
$s(A,B)=(a_{1},\,b_{1},\,a_{2},\,b_{2},\ldots,\,a_{n},\,b_{n}).$
Now we show Theorem 1.4. Similar arguments have appeared in [1, 8] for real numbers and in [19, Section 5.2] for complex numbers.
Proof of Theorem 1.4.
Assume that $\zeta$ is an algebraic number, so $[\mathbb{Q}(i,\zeta):\mathbb{Q}(i)]\geq 3$ (its HCF is not periodic; see, for example, [13, Theorem 4.2]). Let $(W_{n})_{n\geq 1}$, $(U_{n})_{n\geq 1}$, and $(V_{n})_{n\geq 1}$ be the sequences associated to $\mathbf{B}$ as in Lemma 6.1 and define $(w_{n})_{n\geq 1}$, $(u_{n})_{n\geq 1}$, and $(v_{n})_{n\geq 1}$ as above. We may assume that every $u_{n}$ is even. Indeed, suppose that $u_{n}$ is odd for every large $n$. For such $n$, call $x$ be the last letter of $U_{n}$ and $\widehat{U}_{n}:=\operatorname{pref}(U_{n},u_{n}-1)$, so
$W_{n}U_{n}V_{n}U_{n}=W_{n}\,\widehat{U}_{n}x\,V_{n}\widehat{U}_{n}x$
and we can replace $U_{n}$ with $\widehat{U}_{n}$ and $V_{n}$ with $x\,V_{n}$.
When $\liminf_{n\to\infty}w_{n}<\infty$, the sequence $(-2,1+iB_{1},-2,1+iB_{2},\ldots)$ also has finite repetition exponent and its corresponding sequence $(W_{n})_{n\geq 1}$ has a constant subsequence. The transcendence of $\zeta$ then follows from Theorem 6.2. Suppose that $w_{n}\to\infty$ when $n\to\infty$. As noted in [8, p. 1013], we do not lose any generality by assuming $B_{w_{n}}\neq B_{t_{n}}$ for all $n\in\mathbb{N}$. By Example 1, we have $\mathbf{a}\in\mathsf{Ir}$ and the output $\mathbf{b}$ of Algorithm 4.3 is
$\mathbf{b}=(b_{1},b_{2},\ldots)=(-2,iB_{1},2,iB_{2},-2,iB_{3},2,iB_{4},\ldots).$
Note that
(18) $\forall n\in\mathbb{N}\quad\{(b_{1},\ldots,b_{n}),\;(b_{n},\ldots,b_{1})\}\subseteq\mathsf{R}(n).$
For each $n\in\mathbb{N}$, let $A_{n}$, $B_{n}$, $C_{n}$ be the alternating words
[…]
$C_{n}:=\left((-1)^{w_{n}+1}2,\,(-1)^{w_{n}+2}2,\ldots,\,(-1)^{w_{n}+v_{n}}2\right),$
and define
$W_{n}^{\prime}:=s(A_{n},W_{n}),\quad U_{n}^{\prime}:=s(B_{n},U_{n}),\quad V_{n}^{\prime}:=s(C_{n},V_{n}).$
Therefore, since for all $n\in\mathbb{N}$ we have $\mathfrak{F}_{2n}(\operatorname{pref}(\mathbf{b},2n))=\mathfrak{F}$ by Lemma 7.1 below and $u_{n}$ is even, the periodic sequence $Z_{n}:=W_{n}^{\prime}\,U_{n}^{\prime}\,V_{n}^{\prime}\,U_{n}^{\prime}\,V_{n}^{\prime}\,U_{n}^{\prime}\cdots$ belongs to $\overline{\mathsf{R}}$ and $\zeta_{n}:=\overline{\Lambda}(Z_{n})$ is a quadratic irrational over $\mathbb{Q}(i)$. Let $(p_{n})_{n\geq 0}$, $(q_{n})_{n\geq 1}$ correspond to $\mathbf{b}$ as in (5). Given the polynomial
$P_{n}(X):=\begin{vmatrix}q_{2w_{n}-1}&q_{2t_{n}-1}\\ q_{2w_{n}}&q_{2t_{n}}\end{vmatrix}X^{2}-\left(\begin{vmatrix}q_{2w_{n}-1}&p_{2t_{n}-1}\\ q_{2w_{n}}&p_{2t_{n}}\end{vmatrix}+\begin{vmatrix}p_{2w_{n}-1}&q_{2t_{n}-1}\\ p_{2w_{n}}&q_{2t_{n}}\end{vmatrix}\right)X+\begin{vmatrix}p_{2w_{n}-1}&p_{2t_{n}-1}\\ p_{2w_{n}}&p_{2t_{n}}\end{vmatrix},$
one can use the standard proof of Lagrange’s theorem on periodic continued fractions, for example, [13, Theorem 4.2], to check that $\zeta_{n}$ is a solution of $P_{n}$. We are going to use $P_{n}$ to obtain three collections of linear forms and then we will apply Theorem 6.3 to get a contradiction. From $\zeta,\zeta_{n}\in\overline{\mathcal{C}}_{2tn+2u_{n}}(b_{1},\ldots,b_{2tn+2u_{n}})$ and Proposition 3.1.v, we obtain $|\zeta-\zeta_{n}|\leq 2|q_{2t_{n}+2u_{n}}|^{-2}$ (in this estimate we use the triangle inequality as in the proof of Equation (10)). If $\lambda$ is as in Proposition 3.1.vi, we can use parts ii and v of Proposition 3.1 to obtain
$|P_{n}(\zeta)|$ $=|P_{n}(\zeta)-P_{n}(\zeta_{n})|$
$=\left|\zeta-\zeta_{n}\right|\left|\begin{vmatrix}q_{2w_{n}-1}\zeta-p_{2w_{n}-1}&q_{2t_{n}-1}\\ q_{2w_{n}}\zeta-p_{2w_{n}}&q_{2t_{n}}\end{vmatrix}+\begin{vmatrix}q_{2w_{n}-1}&q_{2t_{n}-1}\zeta_{n}-p_{2t_{n}-1}\\ q_{2w_{n}}&q_{2t_{n}}\zeta_{n}-p_{2t_{n}}\end{vmatrix}\right|$
(19) $\leq 2\lambda\left|\zeta-\zeta_{n}\right|\left(\left|\frac{q_{2t_{n}}}{q_{2w_{n}}}\right|+\left|\frac{q_{2w_{n}}}{q_{2t_{n}}}\right|\right)\leq\kappa_{1}\frac{|q_{2t_{n}}|}{|q_{2w_{n}}q^{2}_{2t_{n}+2u_{n}}|}.$
Let $(\mathbf{x}_{n})_{n\geq 1}$, $\mathbf{x}_{n}=(x_{n,1},x_{n,2},x_{n,3},x_{n,4})$, be the sequence in $\mathbb{Z}[i]^{4}$ given by
$\mathbf{x}_{n}:=\left(\begin{vmatrix}q_{2w_{n}-1}&q_{2t_{n}-1}\\ q_{2w_{n}}&q_{2t_{n}}\\ \end{vmatrix},\;\begin{vmatrix}q_{2w_{n}-1}&p_{2t_{n}-1}\\ q_{2w_{n}}&p_{2t_{n}}\\ \end{vmatrix},\;\begin{vmatrix}p_{2w_{n}-1}&q_{2t_{n}-1}\\ p_{2w_{n}}&q_{2t_{n}}\\ \end{vmatrix},\;\begin{vmatrix}p_{2w_{n}-1}&p_{2t_{n}-1}\\ p_{2w_{n}}&p_{2t_{n}}\\ \end{vmatrix}\right).$
Then, $\|\mathbf{x}_{n}\|\leq 2|q_{2w_{n}}q_{2t_{n}}|$ for all $n\in\mathbb{N}$. If $M=\sup\{B_{n}:n\in\mathbb{N}\}$, for all $n\in\mathbb{N}$ we have
$\left|\zeta x_{n,1}-x_{n,2}\right|\leq(M+2)\left|\frac{q_{2w_{n}}}{q_{2t_{n}}}\right|.$
Certainly, by Proposition 3.1. ii, v and Equation (5),
$\left|\zeta x_{n,1}-x_{n,2}\right|$ $=\begin{vmatrix}q_{2w_{n}-1}&\zeta q_{2t_{n}-1}-p_{2t_{n}-1}\\ q_{2w_{n}}&\zeta q_{2t_{n}}-p_{2t_{n}}\\ \end{vmatrix}$
$=\left|q_{2w_{n}-1}(\zeta q_{2t_{n}}-p_{2t_{n}})-q_{2w_{n}}(\zeta q_{2t_{n}-1}-p_{2t_{n}-1})\right|$
$\leq\left|\frac{q_{2w_{n}-1}}{q_{2t_{n}}}\right|+\left|\frac{q_{2w_{n}}}{q_{2t_{n}-1}}\right|\leq(M+2)\left|\frac{q_{2w_{n}}}{q_{2t_{n}}}\right|.$
We can obtain a similar bound for $|\zeta x_{n,1}-x_{n,3}|$. Consider the linear forms $\mathscr{L}_{1,1}$, $\mathscr{L}_{1,2}$, $\mathscr{L}_{1,3}$, $\mathscr{L}_{1,4}$ in the variable $\mathbf{X}=(X_{1},X_{2},X_{3},X_{4})$ and the linear forms $\mathscr{M}_{1,1}$, $\mathscr{M}_{1,2}$, $\mathscr{M}_{1,3}$, $\mathscr{M}_{1,4}$ in the variable $\mathbf{Y}=(Y_{1},Y_{2},Y_{3},Y_{4})$ given by
$\mathscr{L}_{1,1}(\mathbf{X})$ $=\zeta^{2}X_{1}-\zeta(X_{2}+X_{3})+X_{4},$ $\quad\mathscr{M}_{1,1}(\mathbf{Y})=\overline{\mathscr{L}_{1,1}(\overline{\mathbf{Y}})},$
$\mathscr{L}_{1,2}(\mathbf{X})$ $=\zeta X_{1}-X_{2},$ $\quad\mathscr{M}_{1,2}(\mathbf{Y})=\overline{\mathscr{L}_{1,2}(\overline{\mathbf{Y}})},$
$\mathscr{L}_{1,3}(\mathbf{X})$ $=\zeta X_{1}-X_{3},$ $\quad\mathscr{M}_{1,3}(\mathbf{Y})=\overline{\mathscr{L}_{1,3}(\overline{\mathbf{Y}})},$
$\mathscr{L}_{1,4}(\mathbf{X})$ $=X_{1},$ $\quad\mathscr{M}_{1,4}(\mathbf{Y})=\overline{\mathscr{L}_{1,4}(\overline{\mathbf{Y}})}.$
Take $\varepsilon>0$ as in Lemma 6.4. Then, for some positive constants $\kappa_{2},\kappa_{3},\kappa_{4}$ and all $n\in\mathbb{N}$ we have
$\left|\mathscr{L}_{1,1}(\mathbf{x}_{n})\mathscr{L}_{1,2}(\mathbf{x}_{n})\mathscr{L}_{1,3}(\mathbf{x}_{n})\mathscr{L}_{1,4}(\mathbf{x}_{n})\right|$ $=\left|P_{n}(\zeta)\right|\left|\zeta x_{n,1}-x_{n,2}\right|\left|\zeta x_{n,1}-x_{n,3}\right|\left|x_{n,1}\right|$
$\leq\kappa_{2}\left|\frac{q_{2t_{n}}}{q_{2w_{n}}q^{2}_{2t_{n}+2u_{n}}}\frac{q_{2w_{n}}}{q_{2t_{n}}}\frac{q_{2t_{n}}}{q_{2w_{n}}}q_{2w_{n}}q_{2t_{n}}\right|$
$=\kappa_{2}\left|\frac{q_{2t_{n}}^{2}}{q_{2t_{n}+2u_{n}}^{2}}\right|$
$\leq\frac{\kappa_{3}}{\psi^{4u_{n}}}\leq\frac{\kappa_{3}}{|q_{2w_{n}}q_{2t_{n}}|^{\varepsilon}}\leq\frac{\kappa_{4}}{\|\mathbf{x}_{n}\|^{\varepsilon}}.$
In the last line, the first inequality uses Proposition 3.1.vii. The definition of $\mathscr{M}_{1,1}$, $\mathscr{M}_{1,2}$, $\mathscr{M}_{1,3}$, $\mathscr{M}_{1,4}$, tells us that for each $n\in\mathbb{N}$ we have
$\left|\mathscr{M}_{1,1}(\overline{\mathbf{x}_{n}})\mathscr{M}_{1,2}(\overline{\mathbf{x}_{n}})\mathscr{M}_{1,3}(\overline{\mathbf{x}_{n}})\mathscr{M}_{1,4}(\overline{\mathbf{x}_{n}})\right|=\left|\mathscr{L}_{1,1}(\mathbf{x}_{n})\mathscr{L}_{1,2}(\mathbf{x}_{n})\mathscr{L}_{1,3}(\mathbf{x}_{n})\mathscr{L}_{1,4}(\mathbf{x}_{n})\right|.$
Therefore, by Theorem 6.3, there is some $\mathbf{x}=(x_{1},x_{2},x_{3},x_{4})\in\mathbb{Z}[i]^{4}$, $\mathbf{x}\neq(0,0,0,0)$, and an infinite set $\mathcal{N}_{1}\subseteq\mathbb{N}$ such that
(20) $\forall n\in\mathcal{N}_{1}\quad x_{1}x_{n,1}+x_{2}x_{n,2}+x_{3}x_{n,3}+x_{4}x_{n,4}=0.$
Let $(Q_{n})_{n\geq 1}$ and $(R_{n})_{n\geq 1}$ be the sequences given by
$\forall n\in\mathbb{N}\qquad Q_{n}:=\frac{q_{2w_{n}-1}q_{2t_{n}}}{q_{2w_{n}}q_{2t_{n}-1}},\quad R_{n}:=\zeta-\frac{p_{2n}}{q_{2n}}.$
For each $n\in\mathcal{N}_{1}$, we divide Equation (20) by $q_{2w_{n}}q_{2t_{n}-1}$ to obtain
00 0 $=x_{1}(Q_{n}-1)+x_{2}\left(Q_{n}(\zeta-R_{2t_{n}})-(\zeta-R_{2t_{n}-1})\right)+$
$\;+x_{3}\left(Q_{n}(\zeta-R_{2w_{n}-1})-(\zeta-R_{2w_{n}})\right)+$
$\;+x_{4}\left(Q_{n}(\zeta-R_{2w_{n}-1})(\zeta-R_{2t_{n}})-(\zeta-R_{2w_{n}})(\zeta-R_{2t_{n}-1})\right).$
For instance, the coefficient of $x_{2}$ is
$\frac{x_{n,2}}{q_{2w_{n}}q_{2t_{n}-1}}=Q_{n}\frac{p_{2t_{n}}}{q_{2t_{n}}}-\frac{p_{2t_{n}-1}}{q_{2t_{n}-1}}=Q_{n}(\zeta-R_{2tn})-(\zeta-R_{2t_{n}-1}).$
Since $R_{n}\to 0$ as $n\to\infty$, we conclude that
$\lim_{\begin{subarray}{c}n\to\infty\\ n\in\mathcal{N}_{1}\end{subarray}}(Q_{n}-1)\left(x_{1}+(x_{2}+x_{3})\zeta+\zeta^{2}x_{4}\right)=0.$
Let us show that
(21) $x_{1}+(x_{2}+x_{3})\zeta+\zeta^{2}x_{4}=0.$
By Equation (17), Proposition 3.1.ii, and the boundedness of $\mathbf{B}$, we conclude that
$0<\liminf_{n\to\infty}\left|\frac{q_{2w_{n}-1}}{q_{2w_{n}}}\right|<\limsup_{n\to\infty}\left|\frac{q_{2w_{n}-1}}{q_{2w_{n}}}\right|<\infty,$
and the same for $(q_{2t_{n}-1}/q_{2t_{n}})_{n\geq 1}$. Thus, for some infinite subset $\mathcal{N}_{1}^{\prime}\subseteq\mathcal{N}_{1}$ and some constants $a,b\in\mathbb{Z}$, $a\neq b$, we have $a=B_{w_{n}}$ and $b=B_{t_{n}}$ for all $n\in\mathcal{N}_{1}^{\prime}$ and the next limits exist:
$\gamma_{1}:=\lim_{\begin{subarray}{c}n\to\infty\\ n\in\mathcal{N}_{1}^{\prime}\end{subarray}}\frac{q_{2w_{n}}}{q_{2w_{n}-1}},\quad\gamma_{2}:=\lim_{\begin{subarray}{c}n\to\infty\\ n\in\mathcal{N}_{1}^{\prime}\end{subarray}}\frac{q_{2t_{n}}}{q_{2t_{n}-1}}.$
By (18), $\gamma_{1}\in\tau_{ia}\left[\overline{\mathcal{C}}_{1}(-2)\cup\overline{\mathcal{C}}_{1}(2)\right]$ and $\gamma_{2}\in\tau_{ib}\left[\overline{\mathcal{C}}_{1}(-2)\cup\overline{\mathcal{C}}_{1}(2)\right]$ and, since these compact sets are disjoint (see Figure 2), $(Q_{n})_{n\in\mathcal{N}_{1}}$ is bounded away from $1$ for large $n\in\mathcal{N}_{1}$. Hence, (21) holds and, by $[\mathbb{Q}(i,\zeta),\mathbb{Q}(i)]\geq 3$,
$x_{1}=x_{2}+x_{3}=x_{4}=0,\;\text{ so }\;x_{2}=-x_{3}\neq 0.$
As a consequence, by (20), for $n\in\mathcal{N}_{1}$ we have $x_{n,2}=x_{n,3}$, hence
(22) $P_{n}(X)=\begin{vmatrix}q_{2w_{n}-1}&q_{2t_{n}-1}\\ q_{2w_{n}}&q_{2t_{n}}\end{vmatrix}X^{2}-2\begin{vmatrix}q_{2w_{n}-1}&p_{2t_{n}-1}\\ q_{2w_{n}}&p_{2t_{n}}\end{vmatrix}X+\begin{vmatrix}p_{2w_{n}-1}&p_{2t_{n}-1}\\ p_{2w_{n}}&p_{2t_{n}}\end{vmatrix}.$
Let $(\mathbf{v}_{n})_{n\geq 1}$, $\mathbf{v}_{n}=(v_{n,1},v_{n,2},v_{n,3})\in\mathbb{Z}[i]^{3}$, be given by
$\forall n\in\mathbb{N}\quad\mathbf{v}_{n}:=\left(\begin{vmatrix}q_{2w_{n}-1}&q_{2t_{n}-1}\\ q_{2w_{n}}&q_{2t_{n}}\end{vmatrix},\begin{vmatrix}q_{2w_{n}-1}&p_{2t_{n}-1}\\ q_{2w_{n}}&p_{2t_{n}}\end{vmatrix},\begin{vmatrix}p_{2w_{n}-1}&p_{2t_{n}-1}\\ p_{2w_{n}}&p_{2t_{n}}\end{vmatrix}\right),$
so $\|\mathbf{v}_{n}\|_{\infty}\leq 2|q_{2w_{n}}q_{2t_{n}}|$. Consider the linear forms $\mathscr{L}_{2,1}$, $\mathscr{L}_{2,2}$, $\mathscr{L}_{2,3}$ in the variable $\mathbf{X}=(X_{1},X_{2},X_{3})$ and the linear forms $\mathscr{M}_{2,1}$, $\mathscr{M}_{2,2}$, $\mathscr{M}_{2,3}$ in the variable $\mathbf{Y}=(Y_{1},Y_{2},Y_{3})$ given by
$\mathscr{L}_{2,1}(\mathbf{X})$ $=\zeta^{2}X_{1}-2\zeta X_{2}+X_{3},$ $\quad\mathscr{M}_{2,1}(\mathbf{Y})=\overline{\mathscr{L}_{2,1}(\overline{\mathbf{Y}})},$
$\mathscr{L}_{2,2}(\mathbf{X})$ $=\zeta X_{1}-X_{2},$ $\quad\mathscr{M}_{2,2}(\mathbf{Y})=\overline{\mathscr{L}_{2,2}(\overline{\mathbf{Y}}),}$
$\mathscr{L}_{2,3}(\mathbf{X})$ $=X_{1},$ $\quad\mathscr{M}_{2,3}(\mathbf{Y})=\overline{\mathscr{L}_{2,3}(\overline{\mathbf{Y}}).}$
For all $n\in\mathbb{N}$ we have
$\left|\mathscr{M}_{2,1}(\overline{\mathbf{v}_{n}})\mathscr{M}_{2,2}(\overline{\mathbf{v}_{n}})\mathscr{M}_{2,3}(\overline{\mathbf{v}_{n}})\right|=\left|\mathscr{L}_{2,1}(\mathbf{v}_{n})\mathscr{L}_{2,2}(\mathbf{v}_{n})\mathscr{L}_{2,3}(\mathbf{v}_{n})\right|.$
By (19), (22), and parts ii and v of Proposition 3.1, there are some constants $\kappa_{5},\kappa_{6},\kappa_{7}>0$ depending on $\zeta$ such that for all $n\in\mathcal{N}_{1}$ we have
$\left|\mathscr{L}_{2,1}(\mathbf{v}_{n})\mathscr{L}_{2,2}(\mathbf{v}_{n})\mathscr{L}_{2,3}(\mathbf{v}_{n})\right|$ $=|P_{n}(\zeta)(\zeta v_{n,1}-v_{n,2})v_{n,1}|$
$\leq\kappa_{5}\frac{|q_{2t_{n}}|}{|q_{2w_{n}}q^{2}_{2t_{n}+2u_{n}}|}\begin{vmatrix}q_{2w_{n}-1}&q_{2t_{n}-1}\zeta-p_{2t_{n}-1}\\ q_{2w_{n}}&q_{2t_{n}}\zeta-p_{2t_{n}}\end{vmatrix}|q_{2w_{n}}|$
$\leq\kappa_{5}\frac{|q_{2w_{n}}|}{|q_{2t_{n}+2u_{n}}|}\leq\kappa_{5}\frac{|q_{2t_{n}}|}{|q_{2t_{n}+2u_{n}}|}\leq\frac{\kappa_{6}}{|q_{2w_{n}}q_{2t_{n}}|^{\varepsilon}}\leq\frac{\kappa_{7}}{\|\mathbf{v}_{n}\|^{\varepsilon}}.$
The next to last inequality uses Lemma 6.4 and Proposition 3.1.vii. Hence, Theorem 6.3 implies the existence of some infinite set $\mathcal{N}_{2}\subseteq\mathcal{N}_{1}$ and some $\mathbf{r}=(r_{1},r_{2},r_{3})\in\mathbb{Z}[i]^{3}$, $\mathbf{r}\neq(0,0,0)$, such that
$\forall n\in\mathcal{N}_{2}\quad r_{1}v_{n,1}+r_{2}v_{n,2}+r_{3}v_{n,3}=0.$
Dividing by $q_{2w_{n}}q_{2t_{n}-1}$ we obtain
$\lim_{\begin{subarray}{c}n\to\infty\\ n\in\mathcal{N}_{2}\end{subarray}}(Q_{n}-1)(r_{3}\zeta^{2}+r_{2}\zeta+r_{1})=0.$
Since $(Q_{n})_{n\in\mathcal{N}_{1}}$ is bounded away from $1$, we conclude
$r_{3}\zeta^{2}+r_{2}\zeta+r_{1}=0.$
This implies $[\mathbb{Q}(i,\zeta):\mathbb{Q}(i)]\leq 2$, contradicting our hypothesis. Therefore, $\zeta$ is transcendental. ∎
7. Proof of the algorithm
In this section, we demonstrate the validity of Algorithm 4.3, thereby proving Lemma 4.1. We begin by presenting some useful observations of HCF and then we prove Algorithm 4.3.
7.1. Preliminary lemmas
In most of the proofs to come, we use implicitly the equations (2) and (3).
Lemma 7.1.
Take $n,k\in\mathbb{N}$, $(x_{1},\ldots,x_{n})\in\mathscr{D}^{n}$, $(d_{1},\ldots,d_{k})\in\mathscr{D}^{k}$, $m\in\mathbb{Z}$ such that $|m|\geq 2$ and $c\in\{m,im\}$. If $(x_{1},\ldots,x_{n},c,d_{1},\ldots,d_{k})\in\mathsf{R}(n+k+1)$, then
$\mathfrak{F}^{\circ}_{n+k+1}(x_{1},\ldots,x_{n},c,d_{1},\ldots,d_{k})=\mathfrak{F}^{\circ}_{k}(d_{1},\ldots,d_{k}).$
For $k=0$, we have $\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)=\mathfrak{F}^{\circ}(c)$.
Proof.
When $k=0$, the lemma can be shown by checking case by case all the options for $\mathfrak{F}^{\circ}_{n}(x_{1},\ldots,x_{n})$ and $c$. For example, if $\mathfrak{F}^{\circ}_{n}(x_{1},\ldots,x_{n})=\mathfrak{F}^{\circ}_{1}(2)$, then $c=im$ for some $m\in\mathbb{Z}$ with $|m|\geq 2$ or $c=m$ for some $m\in\mathbb{N}_{\geq 2}$ (see Figure 5); in either case, $\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)=\mathfrak{F}^{\circ}(c)$. Assume that $k\geq 1$. If $|c|\geq 3$, then $\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)=\mathfrak{F}^{\circ}$ and, by (8),
$\mathfrak{F}^{\circ}_{n+k+1}(x_{1},\ldots,x_{n},c,d_{1},\ldots,d_{k})$ $=T^{n+k+1}\left[\mathcal{C}^{\circ}_{n+k+1}(x_{1},\ldots,x_{n},c,d_{1},\ldots,d_{k})\right]$
$=T^{k}\left[\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)\cap\mathcal{C}^{\circ}_{k}(d_{1},d_{2},\ldots,d_{k})\right]$
$=T^{k}\left[\mathfrak{F}^{\circ}\cap\mathcal{C}^{\circ}_{k}(d_{1},\ldots,d_{k})\right]$
$=\mathfrak{F}^{\circ}_{k}(d_{1},\ldots,d_{k}).$
When $|c|=2$, we check the $4$ possible forms of $\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)\neq\varnothing$ to conclude that if $(x_{1},\ldots,x_{n},c,d_{1})\in\mathsf{R}(n+2)$ then
$\mathfrak{F}^{\circ}_{n+2}(x_{1},\ldots,x_{n},c,d_{1})=\mathfrak{F}^{\circ}_{1}(d_{1}).$
For instance, if $c=2$, then
$\iota[\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},2)]=\iota\left[\mathfrak{F}^{\circ}_{1}(2)\right]=\iota[\mathfrak{F}^{\circ}]\cap\left\{z\in\mathbb{C}:\operatorname{Re}(z)>\frac{1}{2}\right\},$
so, for all $d\in\mathscr{D}$, we have
$\tau_{d}[\mathfrak{F}^{\circ}]\cap\iota[\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)]=\varnothing\;\text{ or }\;\tau_{d}[\mathfrak{F}^{\circ}]\cap\iota[\mathfrak{F}^{\circ}]$
(cfr. Figures 1 and 4). From here, we argue as in the case $|c|\geq 3$. ∎
Lemma 7.2 follows from the definitions of $\mathsf{mir}_{1}(z)=\overline{z}$, $\mathsf{mir}_{2}(z)=-\overline{z}$, and $S$. Lemma 7.3 is a corollary of Lemma 3.5 and its proof.
Lemma 7.2.
If $m\in\mathbb{Z}$ is such that $|m|\geq 2$ and $a\in\{1+im,m+i\}$, then
$(\mathsf{mir}_{1}\circ S)(a)=(S\circ\mathsf{mir}_{1})(a)\quad\text{ and }\quad(\mathsf{mir}_{2}\circ S)(a)=(S\circ\mathsf{mir}_{2})(a).$
Lemma 7.3.
Take $\mathbf{b}=(b_{n})_{n\geq 1}\in\Omega$. If $j\in\mathbb{N}$ is such that $(b_{1},\ldots,b_{j})\in\mathsf{R}(j)$ but $(b_{1},\ldots,b_{j})\in\mathsf{Ir}(j+1)$, then
$\operatorname{Re}([b_{j};b_{j+1},b_{j+1},\ldots]_{\mathbb{C}})=\frac{1}{2}\quad\text{ or }\quad\operatorname{Im}([b_{j};b_{j+1},b_{j+1},\ldots]_{\mathbb{C}})=\frac{1}{2}.$
Lemma 7.4.
If $\mathsf{mir}\in\mathsf{Mir}$, then $\overline{\Lambda}\circ\mathsf{mir}=\mathsf{mir}\circ\overline{\Lambda}$.
Proof.
The lemma follows from the continuity of $\mathsf{mir}_{1}$ and $\mathsf{mir}_{2}$ from the fact that for every $\mathsf{mir}\in\mathsf{Mir}$ we have
$\mathsf{mir}\left([0;z_{1},\ldots,z_{n}]\right)=[0;\mathsf{mir}(z_{1}),\ldots,\mathsf{mir}(z_{n})]$
for any $z_{1},\ldots,z_{n}\in\mathbb{C}$ such that both sides of the equality are defined. ∎
7.2. Proof of the algorithm
Let $\mathbf{a}=(a_{n})_{n\geq 1}$ be an irregular sequence. Since every cylinder of level $1$ is regular, there is some $j_{0}\in\mathbb{N}$ such that
$(a_{1},\ldots,a_{j_{0}})\in\mathsf{R}(j_{0})\text{ and }(a_{1},\ldots,a_{j_{0}},a_{j_{0}+1})\not\in\mathsf{R}(j_{0}+1).$
Note that $\mathfrak{F}^{\circ}_{j_{0}}(a_{1},\ldots,a_{j_{0}})\neq\mathfrak{F}^{\circ}$, otherwise $(a_{1},\ldots,a_{j_{0}},a_{j_{0}+1})$ would be regular. The open prototype set $\mathfrak{F}^{\circ}_{j_{0}}(a_{1},\ldots,a_{j_{0}})$ can assume twelve forms (see Figure 1); by the proof of Lemma 3.5, we may conclude the existence of some $m\in\mathbb{Z}$ such that
$|m|\geq 2\quad\text{ and }\quad a_{j_{0}+1}\in\left\{1+im,m+i\right\}.$
Let $t=1$ if $a_{j_{0}+1}=m+i$ and $t=2$ if $a_{j_{0}+1}=1+m$. We define
$\mathbf{b}_{1}:=(a_{1},\ldots,a_{j_{0}},S(a_{j_{0}+1}),\mathsf{mir}_{t}(a_{j_{0}+2}),\mathsf{mir}_{t}(a_{j_{0}+3}),\mathsf{mir}_{t}(a_{j_{0}+4}),\ldots).$
Assume that $|m|\geq 3$, then
$\mathfrak{F}^{\circ}_{j_{0}+1}\left(a_{1},\ldots,a_{j_{0}},S(a_{j_{0}+1})\right)=\mathfrak{F}^{\circ}\quad\text{ and }\quad(b_{1}(1),\ldots,b_{1}(j_{0}+1))\in\mathsf{R}(j_{0}+1).$
If $a_{j_{0}+1}=1+im$, by Lemma 7.3 and $[0;a_{j_{0}+2},a_{j_{0}+3},\ldots]\in\mathfrak{F}$, we have
(23) $\operatorname{Re}\left([a_{j_{0}+1};a_{j_{0}+2},a_{j_{0}+3},\ldots]_{\mathbb{C}}\right)=\frac{1}{2}.$
Since $z=1-\overline{z}$ holds whenever $\operatorname{Re}(z)=\frac{1}{2}$, we obtain
$[a_{j_{0}+1};a_{j_{0}+2},a_{j_{0}+3},\ldots]_{\mathbb{C}}$ $=1-\overline{[a_{j_{0}+1};a_{j_{0}+2},a_{j_{0}+3},\ldots]_{\mathbb{C}}}$
$=[1-\overline{a_{j_{0}+1}};\overline{a_{j_{0}+2}},\overline{a_{j_{0}+3}},\ldots]$
$=[S(a_{j_{0}+1});\mathsf{mir}_{2}(a_{j_{0}+1}),\mathsf{mir}_{2}(a_{j_{0}+2}),\mathsf{mir}_{2}(a_{j_{0}+3}),\ldots].$
Therefore,
$\Lambda(\mathbf{a})$ $=[0;a_{1},\ldots,a_{j_{0}},a_{j_{0}+1},a_{j_{0}+2},\ldots]_{\mathbb{C}}$
$=[0;a_{1},a_{2},\ldots,a_{j_{0}},S(a_{j_{0}+1}),\mathsf{mir}_{2}(a_{j_{0}+2}),\mathsf{mir}_{2}(a_{j_{0}+3})\ldots]$
$=[0;b_{1}(1),b_{1}(2),\ldots,b_{1}(j_{0}),b_{1}(j_{0}+1),b_{1}(j_{0}+2),b_{1}(j_{0}+3),\ldots]=\overline{\Lambda}(\mathbf{b}_{1}).$
Similarly, if $a_{j_{0}+1}=m+i$, then Lemma 7.3 and $[0;a_{j_{0}+2},a_{j_{0}+3},\ldots]\in\mathfrak{F}$ give
(24) $\operatorname{Im}\left([a_{j_{0}+1};a_{j_{0}+2},a_{j_{0}+3},\ldots]_{\mathbb{C}}\right)=\frac{1}{2}.$
Then, since $z=\overline{z}+i$ whenever $\operatorname{Im}(z)=\frac{1}{2}$, we have
$[a_{j_{0}+1};a_{j_{0}+2},a_{j_{0}+3},\ldots]_{\mathbb{C}}$ $=i+\overline{[a_{j_{0}+1};a_{j_{0}+2},a_{j_{0}+3},\ldots]_{\mathbb{C}}}$
$=[i+\overline{a_{j_{0}+1}};\overline{a_{j_{0}+2}},\overline{a_{j_{0}+3}},\ldots]$
$=[S\left(a_{j_{0}+1}\right);\mathsf{mir}_{1}(a_{j_{0}+1}),\mathsf{mir}_{1}(a_{j_{0}+2}),\mathsf{mir}_{1}(a_{j_{0}+3}),\ldots].$
We conclude $\Lambda(\mathbf{a})=\overline{\Lambda}(\mathbf{b}_{1})$.
Assume that $|m|=2$. Then, applying $\iota$ on the sets in Lemma 3.5, we can see that $a_{j_{0}+1}\in\{1+2i,1-2i,2+i,-2+i\}$. Suppose that $a_{j_{0}+1}=1+2i$, then we must have either
$\mathfrak{F}^{\circ}_{j_{0}}(a_{1},\ldots,a_{j_{0}})=\mathfrak{F}^{\circ}_{1}(-1-i)\;\text{ or }\;\mathfrak{F}^{\circ}_{j_{0}}(a_{1},\ldots,a_{j_{0}})=\mathfrak{F}^{\circ}_{1}(-2)$
(to see it, apply $\iota$ on every set in Lemma 3.5 and notice that the only ones leading to an irregular cylinder after concatenating $1+2i$ are $\mathfrak{F}^{\circ}_{1}(-1-i)$ and $\mathfrak{F}^{\circ}_{1}(-2)$). In either case, we have
$\mathfrak{F}^{\circ}_{j_{0}+1}\left(a_{1},\ldots,a_{j_{0}},S(a_{j_{0}+1})\right)=\mathfrak{F}^{\circ}_{j_{0}+1}(a_{1},\ldots,a_{j_{0}},2i)=\mathfrak{F}^{\circ}_{1}(2i)\neq\varnothing.$
A similar argument holds when $a_{j_{0}+1}\in\{1-2i,2+i,2-i\}$. Irrespective to what case occurs when $|m|=2$, we may show $\Lambda(\mathbf{a})=\overline{\Lambda}(\mathbf{b}_{1})$ as when $|m|\geq 3$. To sum up, the sequence $\mathbf{b}_{1}$ has the next properties:
  1. i.
    $(b_{1}(1),\ldots,b_{1}(j_{0}+1))\in\mathsf{R}(j_{0}+1)$,
  2. ii.
    $\Lambda(\mathbf{a})=\overline{\Lambda}(\mathbf{b}_{1})$.
Assume that for some $N\in\mathbb{N}$ we have we are able to perform $N$ iterations of the algorithm and the resulting sequence $\mathbf{b}_{N}$ satisfies:
  1. i.
    $(b_{N}(1),\ldots,b_{N}(j_{N-1}+1))\in\mathsf{R}(j_{N-1}+1)$,
  2. ii.
    $\Lambda(\mathbf{a})=\overline{\Lambda}(\mathbf{b}_{N})$.
  3. iii.
    Either
    (25) $\operatorname{Re}\left([a_{j_{N-1}+1};a_{j_{N-1}+2},\ldots]_{\mathbb{C}}\right)=\frac{1}{2}\;\text{ or }\;\operatorname{Im}\left([a_{j_{N-1}+1};a_{j_{N-1}+2},\ldots]_{\mathbb{C}}\right)=\frac{1}{2}.$
Remark 7.5.
By the definition of $\mathbf{b}_{N}$, for each $M\in\{1,\ldots,N\}$ there exists some $\mathsf{mir}\in\mathsf{Mir}$ such that
$\forall n\in\mathbb{N}\quad b_{M}(j_{M-1}+1+n)=\mathsf{mir}\left(b_{M-1}(j_{M-1}+1+n)\right).$
As a consequence, there is some $\mathsf{mir}^{\prime}\in\mathsf{Mir}$ for which
$\forall n\in\mathbb{N}\quad b_{M}(j_{M-1}+1+n)=\mathsf{mir}^{\prime}\left(a_{j_{M-1}+1+n}\right).$
Next, we discuss the $(N+1)$-th iteration. Suppose there is some $j_{N}\in\mathbb{N}$ satisfying
(26) $\left(b_{N}(1),\ldots,b_{N}(j_{N})\right)\in\mathsf{R}(j_{N})\quad\text{ and }\quad\left(b_{N}(1),\ldots,b_{N}(j_{N}+1)\right)\not\in\mathsf{R}(j_{N}+1).$
Let us show that
(27) $\left(b_{N}(1),\ldots,b_{N}(j_{N}),S(b_{N}(j_{N}+1))\right)\in\mathsf{R}(j_{N}+1),$
(28) $\operatorname{Re}\left([a_{j_{N}+1};a_{j_{N}+2},\ldots]_{\mathbb{C}}\right)=\frac{1}{2}\quad\text{ or }\quad\operatorname{Im}\left([a_{j_{N}+1};a_{j_{N}+2},\ldots]_{\mathbb{C}}\right)=\frac{1}{2},\text{ and}$
(29) $\Lambda(\mathbf{a})=\overline{\Lambda}(\mathbf{b}_{N+1}).$
Note that $j_{N}-j_{N-1}\in\mathbb{N}$ by the induction hypothesis. We consider separately the cases $j_{N}-j_{N-1}=1$ and $j_{N}-j_{N-1}\geq 2$.
Case $j_{N}=j_{N-1}+1$. Since $b_{N}(j_{N})=S(b_{N-1}(j_{N}))=S(b_{N-1}(j_{N-1}+1))$, there is some $m\in\mathbb{Z}$, $|m|\geq 2$, such that $b_{N}(j_{N})\in\{m,im,-m,-im\}$. If we had $|m|\geq 3$, by Lemma 7.1 we would have
$\mathfrak{F}^{\circ}_{j_{N}}(b_{N}(1),\ldots,b_{N}(j_{N}))=\mathfrak{F}^{\circ}_{j_{N-1}+1}(b_{N-1}(1),\ldots,S(b_{N-1}(j_{N-1}+1))=\mathfrak{F}^{\circ},$
which implies $(b_{N}(1),\ldots,b_{N}(j_{N}),b_{N}(j_{N}+1))\in\mathsf{R}(j_{N}+1)$, contradicting (26). Therefore, $|m|=2$ and $b_{N}(j_{N})\in\{-2,-2i,2,2i\}$. Let us further assume that $b_{N}(j_{N})=-2$, the other possibilities are treated similarly. Then, we have
$b_{N-1}(j_{N})=b_{N-1}(j_{N-1}+1)\in\{-2+i,-2-i\}.$
Suppose that $b_{N-1}(j_{N})=-2+i$, an analogous argument works for $-2-i$. Under these new assumptions, Condition (26) and Lemma 7.1 imply
(30) $\mathfrak{F}^{\circ}_{j_{N}}\left(b_{N}(1),\ldots,b_{N}(j_{N})\right)=\mathfrak{F}^{\circ}_{1}(-2)\quad\text{ and }\quad\operatorname{Re}\left(b_{N}(j_{N}+1)\right)\geq 1$
(observe that $\mathfrak{F}^{\circ}_{1}(-2)=-\mathfrak{F}^{\circ}_{1}(2)$, $\iota[\mathfrak{F}^{\circ}_{1}(-2)]=-\iota[\mathfrak{F}(2)]$ and see Figure 5). We claim that
(31) $\exists\mathsf{mir}\in\mathsf{Mir}\quad\forall n\in\mathbb{N}\quad b_{N-1}(j_{N-1}+n)=\mathsf{mir}(a_{j_{N-1}+n}).$
Certainly, when $N=1$, we may consider $\mathsf{mir}$ to be the identity map on $\mathbb{C}$. For $N>1$, the induction hypothesis gives $j_{N-1}\geq j_{N-2}+1$, so $j_{N-1}+n\geq j_{N-2}+n+1$ for all $n\in\mathbb{N}$ and (31) follows from Remark 7.5.
Next, we consider every possibility for $\mathsf{mir}$ in (31) separately. Recall that
$-2+i=b_{N-1}(j_{N})=\mathsf{mir}(a_{j_{N}}).$
  • When $\mathsf{mir}$ is the identity map, we have $a_{j_{N}}=-2+i$ and $\operatorname{Re}(a_{j_{N}+1})\geq 1$. Hence, the set $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})\in\mathsf{Ir}(j_{N})$ is of one of the forms depicted on Figure 8 (this can be checked inductively using (3)). We consider each case independently.
    1. (a)
      If $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})=\{[0;a_{j_{N}+1},a_{j_{N}+2},\ldots]_{\mathbb{C}}\}=\{\zeta_{4}\}$, we conclude (28) from $\operatorname{Re}(\zeta_{4}^{-1})=\frac{1}{2}$. Moreover, we have
      $\left((b_{N-1}(j_{N}+k)\right)_{k\geq 0}$ $=(-2+i,a_{j_{N}+1},a_{j_{N}+2},a_{j_{N}+3},\ldots)$
      $=(-2+i,1+2i,-2+i,1+2i,-2+i,\ldots).$
      Then, the definition of $\mathbf{b}_{N}$ and $j_{N}=j_{N-1}+1$ yield
      $\left((b_{N}(j_{N}+k)\right)_{k\geq 0}=(-2,1-2i,-2-i,1-2i,-2-i,1-2i\ldots).$
      Then, $j_{N+1}=j_{N}+1$ by $(-2,1-2i)\in\mathsf{Ir}(2)$ and
      $\left((b_{N+1}(j_{N}+k)\right)_{k\geq 0}=(-2,-2i,2-i,-1-2i,2-i,-1-2i\ldots).$
      We now argue as in the case $N=0$ to conclude
      $\overline{\Lambda}\left(b_{N}(j_{N}),b_{N}(j_{N}+1),\ldots\right)=\overline{\Lambda}\left(b_{N+1}(j_{N}),b_{N+1}(j_{N}+1),\ldots\right)$
      and, since $b_{N}(k)=b_{N+1}(k)$ for all $k\in\{1,\ldots,j_{N}-1\}$, Equation (29) follows.
    2. (b)
      Suppose that $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})$ is the arch of $C(1-i)$ going from $-\zeta_{1}=\frac{1}{2}-i\alpha$ (excluded) to $\zeta_{4}$ (included). Then,
      $\left\{w\in\mathbb{C}:\operatorname{Im}(w)=\frac{1}{2}\right\}\cap\iota[\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})]$ ==$\varnothing$,= ∅ ,
      $\left\{w\in\mathbb{C}:\operatorname{Re}(w)=\frac{1}{2}\right\}\cap\iota[\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})]$ =$=\{\tau_{1+2i}(\zeta_{2})\},$
      so $\zeta_{2}=[0;a_{j_{N+1}},a_{j_{N+2}},\ldots]$ and we argue as in Case (a).
    3. (c)
      When $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})$ is the arch of $C(1)$ from $\overline{\zeta_{4}}$ (excluded) to $\zeta_{4}$ (included), we argue as in case (b).
    4. (d)
      When $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})=[-\frac{1+i}{2},\zeta_{4}]$, we apply $\iota$ on $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})$ to get $a_{j_{N}+1}\in\{-1+i,-1+2i,2i,1+2i\}$ and, since $\operatorname{Re}(a_{j_{N}+1})\geq 1$, we must have $a_{j_{N}+1}=1+2i$. Then,
      $\iota\left[\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})\right]\cap\tau_{a_{j_{N}+1}}[\mathfrak{F}]=\left\{\tau_{a_{j_{N}}+1}(\zeta_{2})\right\}$
      and $\zeta_{2}=[0;a_{j_{N}+1},a_{j_{N}+2},\ldots]_{\mathbb{C}}$. Now, we proceed as in case (a).
    As we will see, the other options for $\mathsf{mir}$ are impossible.
  • If $\mathsf{mir}=\mathsf{mir}_{1}$, then $a_{j_{N}}=-2-i$ but (25) cannot hold.
  • When $\mathsf{mir}=\mathsf{mir}_{2}$, then $a_{j_{N}}=2+i$, $\operatorname{Re}(a_{j_{N}+1})\leq-1$ and, by (25),
    (32) $\operatorname{Im}([a_{j_{N}};a_{j_{N}+1},a_{j_{N}+2},\ldots]=\frac{1}{2}.$
    The possible options for $\mathfrak{F}^{\circ}_{j_{N}}(a_{1},\ldots,a_{j_{N}})$ are depicted in Figure 8.
    1. (a)
      If $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})=\{\zeta_{2}\}=\{[0;a_{j_{N}+1},a_{j_{N}+2},\ldots]\}$, then $[a_{j_{N}};a_{j_{N}+1},\ldots]=-2+i+\zeta_{2}$, contradicting (32).
    2. (b)
      If $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})$ were the arch of $C(-1-i)$ going from $\zeta_{3}$ (included) to $\zeta_{2}$ (included), by (25) we would obtain $\zeta_{3}=[0;a_{j_{N}+1},a_{j_{N}+2},\ldots]_{\mathbb{C}}$, but $a_{j_{N}+1}=2i$ contradicts $\operatorname{Re}(a_{j_{N}+1})\leq-1$.
    3. (c)
      The impossibility of $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})$ being the arch of $C(-1)$ from $\zeta_{2}$ (included) to $\overline{\zeta_{2}}$ (excluded) is shown as in case (b).
    4. (d)
      We cannot have $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})=[\zeta_{3},\frac{1-i}{2})$, because it would imply $a_{j_{N}+1}\in\{2i,1+2i,1+i\}$ which is incompatible with $\operatorname{Re}(a_{j_{N}+1})\leq-1$.
  • When $\mathsf{mir}=\mathsf{mir}_{1}\mathsf{mir}_{2}$, then $a_{j_{N}}=2-i$ and $\operatorname{Re}(a_{{j_{N}}+1})\leq-1$. However, we cannot have (25) if $a_{j_{N}}=2-i$.
Figure 7. Possible forms of $\mathfrak{F}(a_{1},\ldots,a_{j_{N}+1})$ when $\mathsf{mir}$ is the identity map.
Figure 8. Possible forms of $\mathfrak{F}(a_{1},\ldots,a_{j_{N}+1})$ when $\mathsf{mir}=\mathsf{mir}_{2}$.
Case $j_{N}\geq j_{N-1}+2$. By construction, for some $m\in\mathbb{Z}$ with $|m|\geq 2$, we have
$b_{N}(j_{N-1}+1)=S\left(b_{N-1}(j_{N-1}+1)\right)\in\left\{m,im\right\}.$
Then, by Lemma 7.1,
$\mathfrak{F}^{\circ}_{j_{N}}\left(b_{N}(1),\ldots,b_{N}(j_{N})\right)$ $=\mathfrak{F}^{\circ}_{j_{N}-j_{N-1}-1}(b_{N}(j_{N-1}+2),\ldots,b_{N-1}(j_{N}))\neq\varnothing,$
$\mathfrak{F}^{\circ}_{j_{N}+1}\left(b_{N}(1),\ldots,b_{N}(j_{N}+1)\right)$ $=\mathfrak{F}^{\circ}_{j_{N}-j_{N-1}}\left(b_{N}(j_{N-1}+2),\ldots,b_{N}(j_{N}),b_{N}(j_{N}+1)\right)=\varnothing.$
and, by Remark 7.5, we may pick $\mathsf{mir}\in\mathsf{Mir}$ such that
$\forall n\in\mathbb{N}\quad b_{N}(j_{N-1}+1+n)=\mathsf{mir}\left(a_{j_{N-1}+1+n}\right).$
As a consequence-by the choice of $j_{N}$, Proposition 3.4 and Lemma 7.1-we have
$\left(a_{j_{N-1}+2},\ldots,a_{j_{N}}\right)\in\mathsf{R}(j_{N}-j_{N-1}-1)\;\text{ and }\;\left(a_{j_{N-1}+2},\ldots,a_{j_{N}+1}\right)\in\mathsf{Ir}(j_{N}-j_{N-1}-1).$
Applying the case $N=0$ we get
$\left(a_{j_{N-1}+2},\ldots,a_{j_{N}},S(a_{j_{N}+1})\right)\in\mathsf{R}(j_{N}-j_{N-1}-1)$
and (28) follows. We now use Lemma 7.3 to conclude (27).
For $\mathbf{x}=(x_{n})_{n\geq 1}\in\mathscr{D}^{\mathbb{N}}$ write $\mathsf{mir}(\mathbf{x})=(\mathsf{mir}(x_{j}))_{j\geq 1}$. Let $\mathbf{d}$ be the sequence in $\mathscr{D}$ obtained from the first iteration of the algorithm applied on $\mathbf{a}^{\prime}:=(a_{n})_{n\geq j_{N-1}+2}$, then $\Lambda(\mathbf{a}^{\prime})=\overline{\Lambda}(\mathbf{d})$. If $\mathbf{b}^{\prime}:=(b_{N}(n))_{n\geq j_{N-1}+2}$ and $\mathbf{e}:=(b_{N+1}(n))_{n\geq j_{N-1}+2}$, then $\mathsf{mir}(\mathbf{a}^{\prime})=\mathbf{b}^{\prime}$, $\mathsf{mir}(\mathbf{d})=\mathbf{e}$ (see Lemma 7.2) and Lemma 7.4 yields
$\Lambda(\mathsf{mir}(\mathbf{a}^{\prime}))=\mathsf{mir}\left(\Lambda(\mathbf{a}^{\prime})\right)=\mathsf{mir}\left(\overline{\Lambda}(\mathbf{d})\right)=\overline{\Lambda}(\mathsf{mir}(\mathbf{d}))=\overline{\Lambda}(\mathbf{e})$
and (29) follows.
8. Questions for further studies
In recent years, the theory of continued fractions in certain imaginary quadratic number fields has attracted some attention [16, 17, 18]. Although our work uses the symmetries of the HCF process, we believe that a general procedure can be obtained to cover the continued fraction algorithms used in the papers mentioned above. Hence, one of the open questions is to explore:
Question 1.
How can the results of this paper be extended to the continued fraction algorithms in a certain number field such as those studied in [16, 17, 18]?
Thermodynamic formalism studies measures of maximal entropy and equilibrium measures of dynamical systems. This area has interesting interpretations in physics and (fractal) dimension theory (like Bowen’s formula [7]). Thermodynamic formalism for non-finite alphabets is an active area with interesting connections to surface dynamics (see the survey [33]). Unlike the finite-alphabet theory, techniques and results are dependent on the class of systems.
Question 2.
Are there natural equilibrium measures for $(\overline{\mathsf{R}},\sigma)$ and does the variational principle hold?
Another way to represent freedom in a dynamical system is through the space of invariant measures. The richest possible simplex of invariant measures one can attain is the Poulsen simplex. The Poulsen simplex is achieved for finite-alphabet shifts of finite type and some countable alphabet subshifts [22].
Question 3.
What can we say about the simplex of invariant measures of $(\overline{\mathsf{R}},\sigma)$?
Acknowledgements. The authors thank anonymous referees for the careful reading of the manuscript and a number of comments that have improved the quality of this work. We also thank Nikita Shulga for his comments and suggestions
Research by García-Ramos was funded by the Strategic Excellence Initiative program of the Jagiellonian University with grant U1U/W16/NO/01.03. Research by González Robert was partially funded by CONAHCyT through the program Estancias posdoctorales por México. Both González Robert and Hussain were funded by the ARC Discovery Project 200100994.